A Complete Bibliography of the Journal of Number Theory (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

16 July 2020 Version 1.00

Title word cross-reference

(α, β) [173]. (nα) [482]. (p, p) [86]. −2 [651]. 0 [62]. 0, 1 [395]. 1 [62, 98, 521]. 16 [593]. 2 [98, 328, 618, 666, 756]. 24 [591]. 28 [613]. 2^{n} [64]. 2Y^{2} + 1 = Z^{2} [300]. 3 [666]. 32 [147]. 4 [30, 325, 714, 767]. 40 [618]. 4n = 1x + 1y + 1z [112]. 5Y^{2} − 20 = X^{2} [300]. 6 [705]. 6p + 1 [584]. 8 [170, 445, 513, 760]. = Δ + Δ + Δ [463]. > y [405, 435]. ^{2} [256, 274]. ^{3} [256, 274]. A [285, 491, 647]. a^{Ω(n)} [285]. A_{n} [194]. A log |η(z)| [447]. ≈ SL_{2} [602]. ax^{2} + by^{2} + cz^{2} = 0 [202]. b [647, 710]. B_{h}[g] [664]. \binom{(p−1)/2}{(p−1)/4} [495]. mod [10, 712]. mod1 [565]. modp [161, 336, 477]. modp^{2} [495]. modq [73]. bX^{n} + cY^{n} = dZ^{n} [270]. C_{2} × C_{2} [27]. C_{p} [223, 348]. D_{5} [165]. D_{p} [185]. ℓ [248]. η(z) [240]. f(an + b) − f(cn + d) [216]. f(n) [216]. f(X^{p^{2r}} − aX^{p^{r}} − bX) [61]. F_{p^{s}} [61]. F_{q}[x] [533]. G_{m} [520]. J [146, 406, 524]. K [256, 274, 498, 687]. K_{2} [466]. L [9, 25, 65, 286, 294, 308, 318, 361, 426, 512]. L(12, χ) [294]. λ [10, 457, 575]. λ^{−} [599]. ≤ x [435]. ≤ x [405]. LF (2, 2^{n}) [37]. l ≥ 5 [468]. M [564, 706]. M_{11} [453]. M_{12} [453]. Q [96]. Z_{2} [146]. PGSp(4) [694]. SL(2, 3) [343]. SL(2, Z) [52]. SL(3, Z) [344, 490]. SL(3, Z) \ SL(3, R)/SO(3, R) [649, 735]. μ [89, 512]. N [8, 193, 195, 364, 387, 634, 727]. N − p = P_{3} [288]. n > −4 [364]. n ≤ x [285]. O [681]. Ω [675]. P [31, 51, 78, 270, 298, 311, 339, 342, 462, 471, 484, 512, 535, 555, 584, 620, 652, 699, 750, 757, 789]. p ≡ 1(mod16) [201]. p ≡ 3(mod4) [170]. ϕ(x) = k [751]. π [152, 722]. Q [145, 171, 342, 453, 468, 561, 728]. Q(i) [682]. Q(\sqrt{−7}) [600]. Q(\sqrt{−m}) [760]. Q(\sqrt{2p}) [201]. Q(\sqrt{m}) [760]. Q(\sqrt{±p}) [170]. Q(ζ_{p^{n}}) [762]. Q_{p} [223, 348]. s [730]. S(T) [592]. \sqrt{−19} [321]. ∑_{k=1}^{∞}(k^{2k})^{−1}k^{−n} [355]. ∑_{n≤x} A(f(n)) [450]. ∑_{p≤x} A(f(p)) [450]. Υ [463]. X [256, 274]. X^{2} − mY^{2} = −1, −4 [451]. x^{2} = 4q^{a2} + 4q + 1 [553]. x^{3} + 3y^{3} = 2^{n} [186, 323]. x^{3} + y^{3} = D [149]. x^{3} − 3xy^{2} − y^{3} = 1 [284]. X^{n} + aX^{1} + b [529]. X^{p} + Y^{p} = Z^{p} [270]. X_{0}(125) [518]. X_{1}(13) [705]. ξ [198, 312]. Y [256, 274, 285]. y^{2} − D = 2^{k} [244]. y^{2} = 4q^{n} + 4q + 1 [456]. y^{2} = 4 × 3^{a} + 13 [210]. y^{2} = x^{2} − x [682]. Y^{2} = X^{3} − 2^{3} · 19X + 2 · 19^{2} [321]. y^{2} = x^{3} − 4x + 1 [676]. Z [570]. Z[D_{2m}] [116]. Z[Q_{p}] [562]. Z[ζ_{p} + ζ_{p}^{−1}] [662]. Z_{l} [89]. Z_{p} [457, 698, 745, 762]. ζ [247, 728]. ζ(k) [101]. |S(T)| [237].

-adic [51, 298, 311, 339, 342, 361, 462, 471, 484, 512, 555, 652, 750, 757]. -adique [342, 361]. -adiques [342]. -admissible [561]. -Analogue [728]. -Bernoulli [145]. -class [248, 318, 328, 767]. -closed [620]. -color [634]. -density [681]. -dependence [491]. -dimensional [364]. -divisibility [31]. -estimates [675]. -Euler [728]. -Extensions [65, 89, 146, 457, 698, 745, 762]. -fields [146, 406]. -Folgen [395]. -free [687]. -function [198, 728]. -functions [9, 25, 286, 294, 308, 512]. -group [699]. -invariant [512, 575]. -invariants [599]. -isogeny [171]. -isomorphism [171]. -modules [570]. -motives [521]. -polynomials [343]. -primes [666]. -rank [30, 325, 714]. -series [426]. -th [195]. -theory [498]. -tupel [730].

/Galois [712].

1983 [254].

= [8, 256, 274].

abc [688]. abc-conjecture [688]. Abel [778]. abelian [86, 143, 159, 204, 291, 294, 367, 377, 520, 582, 583, 599, 637, 644, 671, 699, 741]. abélien [86]. absolut [417]. absolut-algebraischen [417]. absolute [194]. absolutely [417]. Abzählreime [560]. Accélération [636]. achten [251]. Addendum [91]. Addition [582, 583, 727, 743]. Additive [83, 213, 216, 236, 243, 269, 458, 505, 625, 642, 755, 761, 770, 790]. adélique [447]. adic [51, 298, 311, 339, 342, 361, 447, 462, 471, 484, 512, 555, 652, 750, 757]. adique [342, 361]. adiques [342]. admissible [561]. Alain [348]. Algebraic [45, 47, 98, 111, 177, 233, 268, 292, 309, 318, 340, 342, 361, 362, 399, 406, 417, 420, 424, 437, 443, 458, 462, 469, 472, 498, 528, 534, 549, 565, 610, 611, 665, 691, 758]. algebraischen [47, 98, 417]. algebras [454, 568, 695, 769]. algébriques [342, 361]. algorithm [22, 167, 365, 366, 533, 546, 547, 586, 604]. Algorithms [571]. allied [97]. almost [55, 579]. alternating [718, 768]. among [195, 527]. analog [240]. analogous [447]. analogue [291, 326, 330, 728]. analogues [447]. anneau [468, 682]. anneaux [624]. Announcement [369, 380, 389]. answer [173]. answers [576]. any [251, 658]. Apery [29, 148, 205, 394, 525]. Application [182, 328, 419, 446, 491, 553, 652]. applications [1, 347, 444, 717]. approach [100, 687]. approximable [322]. Approximants [504]. Approximation [47, 68, 208, 261, 312, 364, 483, 484, 497, 616, 667]. Approximationen [138]. approximations [138, 152, 346]. après [738]. April [129, 206, 279, 356]. arbitrary [612, 775]. Archimedean [199, 757]. arising [104]. Arithmetic [34, 49, 83, 121, 213, 216, 307, 376, 407, 567, 601, 622, 662, 715, 721, 725, 729, 749, 759, 761, 777]. arithmetical [6, 160, 675]. arithmétiques [725]. Artin [150, 207, 291, 437]. assigned [74]. associated [25, 35, 122, 392, 441, 459]. associées [122, 441]. asymmetric [138, 616]. asymmetrische [138]. Asymptotic [26, 126, 164, 248, 378, 433, 625, 722]. attached [341, 512]. August [43, 92, 154, 227, 305, 382]. Author [69, 117, 153, 190, 226, 265, 304, 349, 381, 411, 438, 474, 508, 542, 572, 606, 638, 668, 700, 732, 763, 794]. automorphic [626]. Automorphisms [78]. aux [122, 725, 736]. average [160, 247].

badly [322]. Barban [111]. Base [602, 694, 710]. basen [560]. Bases [26, 249, 251, 378, 467, 560, 625, 633, 641]. basis [12, 571]. behavior [40, 248]. behaviour [601]. Beiträge [74]. beliebigem [251]. belonging [515]. Berechnung [255]. Bernoulli [31, 60, 123, 145, 189, 200, 326, 515, 551, 636, 781]. beschränkter [302]. Beschränkung [203]. Bessel [579]. better [409]. between [144, 232, 285, 424, 494, 782]. Beurling [681]. Bhaskaran [347]. Binäre [432]. binary [12, 99, 174, 188, 196, 209, 309, 331, 362, 408, 432]. binomial [172, 182, 353, 388, 603, 766]. binomials [45]. biquadratic [181]. Birch [336]. Bleicher [485, 686]. Bloch [759]. Board [3, 71, 119, 155, 192, 228, 267, 306, 351, 383, 413, 440, 476, 510, 544, 574, 608, 640, 670, 702, 734, 765]. bodies [81, 417, 464, 705]. body [84, 86, 251, 692]. Bohl [415]. Bohr [257]. bounded [216]. bounds [31, 409, 747]. Boyd [220]. branching [203]. broken [560]. Bundschuh [496].

calcul [636]. calculate [604]. calculation [255, 636]. Caliber [692]. Calibre [692]. cancellation [557, 695]. Canonical [354]. Cantor [278, 779]. capitulation [87]. caractères [581]. Carl [379]. Carlitz [145, 781, 784]. case [240, 339, 359]. Catalan [499]. Cauchy [338]. central [717]. Certain [6, 17, 25, 33, 204, 213, 287, 311, 318, 388, 398, 407, 443, 467, 469, 472, 515, 558, 619, 767, 773, 782]. certaines [447, 467, 558]. chain [360]. chains [522]. change [602, 694]. Character [233, 303, 321, 328, 448, 661]. characteristic [78, 591, 612, 666, 756]. characteristics [417]. characterization [505, 620]. Characters [17, 44, 197, 341, 515, 581, 793]. charakteristik [417]. Chowla [156, 250, 330, 357, 371, 401]. circle [646]. circuit [12]. Class [19, 30, 72, 102, 131, 160, 169– 171, 174, 201, 225, 248, 259, 309, 318, 325, 328, 372, 386, 396, 403, 408, 414, 426, 430, 445, 478, 546, 593, 598, 612, 621, 706, 713, 714, 741, 760, 767, 773]. class-number [430]. classes [52, 171, 230, 287, 397, 682]. Classical [626]. classification [46, 125]. closed [620]. CM [65, 739, 747]. CM-fields [65]. Codecà [675]. codes [210, 749]. codimension [759]. coding [553]. coefficient [301, 419]. coefficients [172, 182, 247, 353, 388, 419, 603, 766]. Cofinite [378]. Cohomological [521]. cohomology [344, 420]. color [634]. colored [782]. combinations [477]. Combinatorial [635]. coming [712]. common [63, 388]. commutative [569]. comparative [452]. complete [323]. completely [73]. Completeness [109]. completion [243, 642]. complex [159, 294, 312, 319, 444, 512, 522, 593, 598, 600, 745, 759, 786]. complex-multiplication [786]. composite [89, 137]. Composition [188]. computation [127, 546]. Computations [344]. Computing [373]. concerning [220, 229, 362]. condition [517]. conductors [712]. configurations [754]. Congruence [14, 29, 205, 344, 525, 663]. Congruences [148, 170, 200, 213, 367, 372, 394, 445, 593, 645, 659, 760, 766]. conics [614]. conjecture [8, 97, 150, 207, 224, 291, 330, 336, 357, 371, 401, 437, 485, 487, 555, 580, 586, 688, 707, 723, 726, 737, 762]. conjugacy [52, 287, 396]. conjugate [458]. connecting [458]. connexes [221]. consécutifs [725]. consecutive [144, 327, 494, 725]. consequences [452]. consist [121]. constant [653]. constants [68, 364]. constraints [72]. constructing [249]. Construction [1, 77, 343, 347, 507]. constructive [310]. containing [397, 622]. Continued [77, 110, 133, 134, 241, 257, 275, 346, 433, 442, 451, 471, 472, 488, 591, 684, 779]. Continuous [566]. contributions [13, 74]. convergents [241]. convex [84]. convolution [83]. convolutions [675]. coprimality [376]. coprime [256]. Coprimely [609]. core [203]. Corps [61, 81, 86, 342, 419, 464, 513, 682, 692, 705]. Corput [222, 415, 677]. correction [465]. Corrigenda [214, 289]. Corrigendum [57, 334]. Cotangent [295]. Counterexample [465]. counting [418, 560]. courbe [518, 705]. courbes [81]. coverings [420]. Critères [419]. criteria [419]. criterion [514, 688]. critical [198, 234, 746]. cube [320, 587]. cube-like [587]. cubic [104, 150, 274, 321, 363, 430, 464, 581, 758]. cubique [581]. cubiques [464]. curve [263, 321, 518, 705]. curves [81, 96, 171, 183, 261, 274, 386, 444, 512, 600, 614, 697, 786]. cuspidal [344]. Cycles [88]. Cyclic [13, 49, 299, 430, 468, 637, 705, 749]. cycliques [468, 705]. Cyclotomic [65, 102, 131, 143, 146, 245, 368, 403, 429, 467, 632, 633, 750]. cyclotomiques [467]. cyclotomy [124].

dans [419, 467, 698]. Darstellung [66, 135]. Davenport [111]. December [179, 253, 335, 400]. decimal [90, 97]. decomposable [107, 136]. decomposition [197, 432, 502]. Dedekind [5, 40, 67, 151, 158, 247, 255, 295, 341, 493, 578, 628]. dedekindscher [255]. Dedication [4]. Deficient [520]. defined [332]. definite [35, 135, 309]. Degenerate [530]. degré [468, 513, 705]. degree [64, 127, 236, 251, 468, 506, 513, 534, 618, 666, 705]. degrees [703]. degrés [703]. delta [754]. delta-lambda [754]. denominator [551]. Denominators [501, 591, 630, 631]. Dense [565]. densité [184, 436]. densities [33, 556, 768]. Density [53, 184, 230, 328, 436, 681, 693]. d’entiers [184, 624]. dependence [491]. d’équations [419]. derivatives [198, 234]. determinant [162, 296]. determination [96, 495, 773, 784]. determine [772]. determining [30]. d’Euler [262]. deux [441]. developments [560]. d’exponentielles [703]. diagonal [51, 150, 645]. diagonals [549]. diameters [406]. diaphony [677]. Dickson [696]. Difference [36, 144, 193, 492, 494, 613, 710]. difference-differential [492]. differences [88, 276, 761]. differential [492]. differentials [78]. differents [611]. digit [293, 785]. digital [425, 710, 776]. digits [278, 392, 423, 433, 488]. Dihedral [152]. dilated [565]. dilogarithm [340]. dimension [498, 635]. dimensional [364, 654, 685]. dimensions [779]. diophantic [47]. diophantine [62, 68, 104, 113, 138, 168, 186, 208, 244, 268, 273, 284, 300, 364, 444, 456, 478, 497, 553, 571, 605, 616, 665, 667]. diophantische [47, 138]. Direct [259]. Dirichlet [73, 74, 122, 280, 308, 313, 341, 441]. Dirichletreihen [74]. Discrepancies [176]. discrepancy [82, 395, 482, 550, 674]. discrete [604]. discriminant [345, 393]. Discriminants [127, 445, 513, 612]. disjoint [238, 276, 641]. diskrepanz [395]. dispersion [674]. distinct [59]. distinction [424]. distributed [488]. distribution [38, 41, 175, 176, 213, 237, 247, 271, 278, 346, 363, 480, 532, 559, 628, 647, 730, 749, 761, 776]. dividers [725]. diviseurs [725]. divisibility [19, 31, 619]. divisible [445]. division [321, 644, 682, 769]. divisor [28, 132, 164, 387, 388, 436, 470, 585, 629, 646, 715]. divisorielle [184, 436]. Divisors [241, 327, 353, 387, 516, 647, 716, 777]. domain [587]. Drinfeld [393]. dualité [736]. duality [736, 759]. d’un [86, 342, 419, 692]. d’une [184, 581, 705]. durch [66, 135]. dyadic [105]. Dyer [336]. Dynamical [792].

each [171]. easy [211]. editor [178, 333, 473, 588, 731]. Editorial [3, 71, 119, 155, 192, 228, 267, 306, 351, 383, 413, 440, 476, 510, 544, 574, 608, 640, 670, 702, 734, 765]. Effective [6, 115, 320, 409]. Egyptian [501, 630, 631]. eighth [251, 613]. Einbettungsprobleme [203]. Einheiten [116]. Einheitenbasen [251]. Einheitengruppen [98]. Eisenstein [598, 685]. Eisenstein-series [685]. Eisensteinian [208]. Elementary [67, 158, 416, 582, 583, 687]. elgenvectors [21]. Elliptic [81, 171, 183, 263, 352, 443, 444, 512, 514, 541, 578, 600, 614, 713, 786]. elliptiques [81, 541]. embedding [203, 375, 698, 717]. ensemble [658]. ensembles [436]. entiers [122, 221, 468, 682, 740]. enumeration [162, 540]. Epstein [40]. equal [100, 397, 480]. equation [18, 112, 113, 186, 202, 210, 235, 244, 270, 272, 284, 315, 456, 490, 499, 553, 595, 711, 729, 751, 788]. Equations [104, 194, 236, 268, 284, 300, 311, 324, 419, 427, 444, 451, 455, 478, 492, 530, 571, 605, 652, 673, 704]. equirepartition [221]. equivalence [136, 209, 258, 362, 590]. equivalent [522]. Erdos [434, 485, 576, 686, 723]. Ernst [254]. Erratum [23, 708]. Erzeugenden [98]. Escassut [348]. essentially [337]. estimate [328]. Estimates [16, 68, 126, 364, 410, 675]. estimating [575]. Estimations [410]. eta [536]. Étude [222]. Euclidean [81]. euclidiens [81]. Euler [108, 163, 262, 728, 751]. Eulerian [281]. Evaluation [321, 416, 448, 581]. even [430, 618]. Every [613]. Exact [625]. example [11]. Examples [561, 618]. exceeding [635]. exceeds [598]. Exceptional [81, 493, 548]. exceptionnelles [81]. expansion [180, 393, 750]. expansions [442, 462, 471, 591, 757, 785]. Explicit [308, 375, 490]. explicites [375]. Exponential [25, 571, 659, 704, 791]. Exponential-polynomial [704]. exponentials [703]. exponentielles [704]. exponentielles-polynômes [704]. expressible [337]. expressions [591]. Extension [15, 187, 338, 557]. Extensions [13, 64, 65, 89, 142, 146, 291, 372, 457, 467, 468, 507, 520, 610, 611, 620, 644, 671, 698, 745, 762, 767]. extremal [39, 470, 618].

facteur [740]. factor [63, 532, 740, 790]. factorisatio [229]. factorisation [61]. Factorization [61, 281, 559, 783]. factors [172, 260, 285, 405, 418, 435, 516, 559, 683, 710]. Fahnentransitive [417]. families [66, 455, 703]. famille [705]. familles [703]. family [165, 521, 705]. February [2, 70, 118, 191, 254, 266, 350, 421, 519, 617, 709]. Fermat [163, 182, 189, 225, 235, 270, 315, 359, 662, 697]. Fibonacci [423]. field [1, 30, 61, 105, 197, 208, 272, 332, 339, 342, 345, 347, 363, 372, 386, 402, 419, 429, 612, 621, 665, 682, 714, 741]. Fields [19, 27, 42, 45, 65, 75, 98, 102, 104, 111, 114, 127, 131, 143, 146, 159, 177, 181, 199, 219, 233, 239, 248, 251, 255, 292, 294, 299, 301, 303, 309, 315, 318, 324, 325, 340, 372, 386, 396, 397, 399, 403, 406, 426, 430, 437, 445, 455, 457, 465, 466, 498, 513, 514, 568, 593, 598, 599, 604, 610, 611, 620, 623, 632, 633, 666, 671, 672, 684, 695– 697, 713, 724, 739, 745, 748, 750, 757, 758, 760, 767, 769, 773, 775]. fifth [100]. Fifty [130]. fini [61]. Finite [61, 114, 203, 219, 301, 324, 540, 610–612, 623, 666, 672, 696, 697, 724, 743]. finitely [570]. first [293, 359]. Flag [417]. Flag-transitive [417]. Folgen [395]. fonction [262]. fonctions [447, 541, 725]. form [272, 427, 490, 503, 530, 587]. formal [301, 326]. formed [82]. Formen [66, 135, 432]. forms [35, 40, 46, 51, 55, 66, 74, 99, 103, 107, 120, 136, 150, 174, 188, 196, 209, 218, 224, 239, 258, 297, 298, 331, 352, 362, 404, 408, 432, 446, 454, 459, 545, 566, 590, 626, 651, 656, 665, 712, 724, 768, 771, 789]. formula [35, 158, 250, 263, 423, 490, 511, 649, 722, 735, 741]. formulae [280]. formulas [164, 358]. four [685]. four-dimensional [685]. Fourier [257]. fourth [251]. fraction [77, 241, 346, 433, 442, 471, 591, 779]. Fractional [132]. fractions [90, 110, 133, 134, 257, 275, 451, 472, 488, 501, 630, 631, 684]. Fredholm [134]. Free [39, 308, 391, 405, 435, 455, 687, 758]. Freiman [576]. French [61, 81, 86, 122, 184, 221, 222, 262, 342, 361, 375, 401, 410, 415, 419, 436, 441, 447, 464, 467, 468, 504, 513, 518, 524, 541, 558, 581, 636, 658, 682, 692, 698, 703– 705, 725, 736, 740]. Frobenius [506, 782]. Führer [44]. full [107, 136]. function [7, 10, 28, 56, 83, 102, 126, 131, 144, 198, 234, 237, 246, 262, 272, 309, 313, 358, 393, 418, 426, 429, 592, 596, 604, 695, 728, 746]. functions [6, 9, 25, 35, 40, 85, 132, 160, 164, 204, 213, 216, 232, 247, 255, 269, 286, 294, 308, 319, 376, 384, 407, 428, 443, 447, 459, 470, 505, 512, 528, 535, 541, 626, 716, 725, 755, 761, 790]. Fundamental [181, 201]. Funktionenkörper [272]. Further [200, 295, 402].

G [254]. g.c.d [564]. Galois [42, 64, 75, 86, 165, 185, 194, 339, 373, 420, 453, 506, 507, 529, 698, 712, 713, 718]. Galoisfeldern [42, 75]. Galoisgruppen [453]. galoisienne [86, 624]. galoisiens [698]. games [173, 322]. gaps [556]. Gauss [539, 663, 678]. Gaussian [44, 76, 448, 777]. Gaußschen [44]. gebrochenen [560]. general [270, 452, 585, 760]. Generalisation [663]. Generalisations [650]. generalisées [222, 415]. Generalization [50, 84, 295, 329, 365, 366, 533, 546, 742]. Generalizations [115, 163]. Generalized [158, 164, 222, 225, 255, 269, 341, 415, 515, 547, 551, 681, 718, 755, 782]. generalizing [182]. generated [570, 644]. generating [419]. Génération [682]. generators [98]. génératrice [419]. genus [1, 120, 282, 299, 347, 402, 465, 557, 590]. geodesics [655]. geometric [100, 220]. géométrie [736]. geometry [264, 424, 479, 736]. Germain [584]. German [41, 42, 44, 47, 66, 74, 75, 98, 116, 138, 203, 251, 255, 272, 302, 325, 374, 395, 417, 432, 453, 480, 560]. given [63, 204, 283, 564]. Gleichverteilung [480]. global [239, 417, 692, 695, 748, 769]. globalen [417]. Goldbach [20, 217, 666]. Grades [251]. Graham [580]. Grahm [707]. grand [740]. Grands [703]. graph [221]. Graphes [221]. Great [703]. greatest [388]. grössencharacter [263]. Größencharaktere [44]. ground [142]. Group [30, 72, 86, 143, 165, 185, 194, 252, 277, 287, 314, 325, 332, 342, 391, 396, 538, 693, 699, 714]. Group-theoretic [72]. groupe [86, 342]. Groupes [464]. groups [98, 105, 131, 204, 230, 252, 259, 283, 287, 314, 318, 367, 373, 377, 417, 420, 426, 453, 464, 506, 507, 529, 534, 538, 540, 561, 582, 583, 610, 614, 637, 691, 718, 748]. Gruppen [417].

H. [330, 357, 371]. Halberstam [111]. half [446]. Halton [677]. Hardy [537]. Hartogs [199]. Hausdorff [779]. having [532, 695]. Hecke [5, 775]. Heilbronn [307]. hereditarily [36]. higher [11]. highest [63]. Hilbert [240, 567, 621]. holomorphic [78]. homogeneous [224, 297, 486, 752]. Hsia [125]. hybrid [552]. Hyperbolic [490, 685, 788]. hyperbolics [490]. hyperelliptic [614]. hypergeometric [448, 626]. hypergéométriques [558]. hypergomometric [558]. hypothesis [452, 707].

ideal [30, 363, 368, 397, 714]. ideals [16, 249, 397]. identically [488].

Identities [101, 151, 211, 634, 781]. identity [5, 793]. ihre [203]. II [38, 123, 133, 137, 164, 218, 234, 236, 297, 357, 366, 391, 408, 481, 482, 506, 538, 557, 583, 631, 704, 716, 778]. III [169, 317, 667]. imaginaires [513]. imaginary [19, 167, 513, 514, 599]. impair [682]. impossibility [270]. including [127]. Indecomposable [141]. indefinite [174, 196, 297, 404, 408, 459]. Indépendance [342, 361]. independence [342, 361, 443, 469]. independent [488]. indeterminates [441]. indéterminées [441]. index [69, 117, 153, 190, 226, 265, 304, 349, 368, 381, 411, 422, 438, 474, 508, 542, 572, 606, 638, 668, 700, 732, 741, 763, 794]. Indices [717]. inequalities [103, 168, 218, 273, 302, 665]. inequality [345, 577, 683]. infinite [392, 622, 706]. inhomogeneous [651]. inner [340]. Integer [433, 613, 676, 683, 706, 710, 772]. integer-valued [772]. Integers [51, 88, 105, 122, 167, 184, 221, 243, 249, 268, 278, 285, 323, 327, 337, 378, 405, 406, 418, 435, 468, 516, 532, 600, 615, 660, 681, 682, 691, 727, 740, 777, 790]. Integral [55, 252, 256, 287, 314, 397, 483, 503, 538, 587, 613, 756, 771]. integrales [441, 558]. integrals [441, 558]. interval [687]. intervals [28, 53, 532]. invariant [89, 512, 524, 575]. invariants [65, 457, 487, 527, 599, 604, 739]. Involutive [446]. involving [126, 158, 200, 423, 793]. irrational [82, 90, 133]. Irrationalité [558]. irrationality [329, 496, 526, 558]. irrationals [442, 522]. Irreducibility [169, 515, 517, 567, 579, 612]. irreducible [36, 339, 591, 683]. irregular [429]. irregularity [422]. isogenies [49]. isogeny [171]. isomorphic [194, 444]. Isomorphism [142, 171]. iterated [83]. Iwasawa [89, 487, 527, 575, 599, 739].

Jackson [224]. Jacobi [338, 726]. Jacobian [518]. jacobienne [518]. Jacobsthal [416]. January [412, 509, 607, 701]. joint [776]. Jordan [55]. July [254, 461, 563, 657, 753]. June [139, 215, 290, 370, 449, 554, 648, 744].

Kennzeichnung [98]. Kern [203]. Khintchine [653]. Klassengruppe [325]. Knopp [67]. known [14]. Körpern [417]. Krishnamurthy [97]. Kronecker [79]. Kummer [514]. Kurokawa [694].

Lagrange [546, 547]. Laguerre [718]. lambda [754]. Lambert [158]. Large [236, 418, 516, 532, 615, 661, 740]. Largest [238, 790]. last [182, 189, 359, 662]. lattice [187, 571]. lattices [125, 162, 309, 484, 618, 654]. Laurent [264]. law [75, 655]. laws [11, 432]. leaders [44]. Least [34, 177, 414]. Legendre [41, 202]. Legendresymbole [41]. Lehmer [79]. lemma [415, 678]. Lemme [415]. Length [501, 547, 631, 658]. Leopoldt [541, 555, 586]. Letter [178, 333, 473, 588, 731]. level [769]. Lévy [693]. lié [736]. liées [725]. liés [524, 705]. lift [738]. like [360, 587]. limit [385]. limited [302]. line [9, 198, 234, 746]. linéaire [342]. linéaires [704]. Linear [157, 342, 352, 417, 428, 477, 481, 491, 704, 785]. lineare [417]. Linearly [324]. linked [524, 705]. Liouville [10]. List [10, 95]. Local [105, 386, 419, 611, 756, 768, 769, 771]. localizers [354]. locally [407]. logarithmes [342, 504]. logarithms [342, 352, 504]. longueur [658]. Lösbarkeit [203]. Lösungsmenge [302]. Low [666, 674]. Low-discrepancy [674]. low-dispersion [674]. Lower [747]. Lubin [301]. Ludwig [379]. Lüroth [653].

M. [347, 485]. Mahler [329, 496, 742]. Maillet [296]. main [487]. mappings [570]. March [431, 531, 627, 720]. marking [98]. Markoff [32, 360]. Markoff-like [360]. mathematician [130]. Mathieu [453]. Mathieugruppen [453]. matrices [183, 483, 523]. Matrix [21, 673]. matroids [12]. Maximal [238, 649]. May [24, 80, 439, 543, 639, 733]. mean [237, 746]. mean-square [746]. means [220]. measure [443]. measures [719]. Mellin [541]. memory [93]. meromorphic [528]. metabelian [538]. metacyclic [252, 287, 314]. method [575, 652]. methods [219]. Metric [98, 667]. Metrische [98]. minima [40, 196, 404, 424]. Minimal [26, 167, 406, 735]. minimum [127]. Minkowski [8, 84, 232, 643]. missing [536]. mittels [255]. mixed [168, 273]. Models [159]. Modifying [643]. modulaire [518, 524, 705]. modular [96, 240, 396, 446, 518, 524, 705, 712]. module [64, 282, 420, 557]. Modules [142, 180, 393, 530, 570, 698]. modulo [38, 170, 175, 535, 593, 706, 727, 760, 789]. Moebius [126]. moments [269, 755]. monogénéité [468]. monomials [791]. monotonic [505]. motives [521]. Multi [59]. Multi-rowed [59]. multiples [82, 436, 730]. multiplication [159, 444, 512, 600, 745, 786]. Multiplicative [48, 85, 204, 458, 552, 716]. multipliers [536]. Müntz [480].

N [485]. narrow [714]. natural [26, 377, 693]. nearest [433, 488]. Nearly [115]. négatifs [122]. negative [122]. neighbourhood [385]. Néron [384]. Nevanlinna [409]. nichtlineare [47]. nilpotent [637]. no [418, 532, 751]. nombre [740]. nombres [342, 361, 513, 636, 736]. Non [105, 177, 199, 224, 297, 468, 713, 757]. non-dyadic [105]. non-homogeneous [224, 297]. non-residues [177]. Nonarchimedean [384]. noneffective [6]. nonlinear [47]. nonmaximal [354]. nonnegative [83]. nonstandard [49, 567]. Nonuniformity [468]. nonvanishing [486, 752]. Norm [272, 302, 427, 503, 505, 530, 587, 748]. norm-form [503, 587]. Normal [45, 53, 77, 467, 507, 658]. normales [467]. normality [502]. Normformgleichung [272]. Normformungleichungen [302]. norms [397, 539, 756, 771]. note [34, 60, 125, 235, 318, 362, 422, 429, 437, 487, 499, 526, 621, 656, 678]. notes [14]. November [58, 106, 500, 597, 690, 787]. num [463]. Number [19, 45, 52, 77, 82, 94, 98, 102, 104, 111, 127, 171, 172, 177, 181, 197, 201, 204, 212, 233, 248, 251, 255, 256, 283, 288, 292, 318, 325, 328, 342, 345, 363, 387, 396, 397, 399, 405, 414, 430, 435, 437, 465, 466, 516, 559, 560, 587, 598, 620, 625, 632, 633, 636, 647, 654, 665, 671, 680, 721, 736, 740, 741, 758, 773, 775]. Number-theoretic [654]. numbers [26, 29, 31, 47, 48, 60, 123, 133, 134, 137, 141, 145, 148, 170, 174, 176, 189, 200, 205, 225, 248, 260, 264, 286, 293, 309, 322, 326, 342, 361–363, 377, 385, 394, 403, 408, 424, 445, 458, 469, 478, 479, 484, 502, 513, 523, 525, 526, 535, 546, 551, 565, 566, 593, 610, 659, 667, 687, 693, 713, 728, 730, 760, 781, 792]. numerorum [229].

Obituary [128]. obstruction [317]. October [166, 242, 316, 390, 489, 589, 679, 774]. odd [48, 236, 682]. often [387]. omega [83, 629]. One [38, 103, 162, 195, 218, 337, 598, 635, 759]. One-sided [103, 218]. only [121, 337, 377]. operators [5]. Oppenheim [229]. order [157, 160, 204, 223, 247, 283, 303, 348, 625]. Orders [141, 354, 377, 695, 783]. ordinary [782]. orthogonal [105, 339]. Oskar [140]. other [757].

P. [485]. packed [609]. Padé [504]. Pages [2, 24, 43, 58, 70, 80, 92, 106, 118, 129, 139, 154, 166, 179, 191, 206, 215, 227, 242, 253, 266, 279, 290, 305, 316, 335, 350, 356, 370, 382, 390, 400, 412, 421, 431, 439, 449, 461, 475, 489, 500, 509, 519, 531, 543, 554, 563, 573, 589, 597, 607, 617, 627, 639, 648, 657, 669, 679, 690, 701, 709, 720, 733, 744, 753, 764, 774, 787]. Pairs [236, 584]. pairwise [238]. Paolo [675]. parabolic [649, 735]. parameterized [251]. parametric [165, 181]. parametrisierte [251]. parametrized [263, 430]. part [132]. partial [251]. partition [15, 535, 780]. Partitionings [120]. Partitions [59, 238, 634, 641, 778, 782]. parts [59, 780]. pattern [10]. Patterns [231]. Paul [93–95]. Pell [451, 490, 788]. Pellian [18]. Perfect [48, 115, 157]. period [547]. Periodic [161, 180]. Periodicities [535, 706]. Periodicity [471]. periods [97]. Permutation [193, 292]. Perron [140]. perturbed [109, 660]. Petersson [67]. Petites [262]. Petits [513]. phenomenon [293]. PHKA [463]. Pisot [385]. Pjateckiĭ [212]. plane [446]. plongement [375]. plongements [698]. point [385]. points [96, 321, 503, 520, 587, 600, 644, 676, 682]. polygons [261]. polylogarithm [399]. polynôme [122, 441]. polynômes [61, 410, 704]. Polynomial [38, 73, 108, 122, 195, 219, 239, 245, 249, 410, 441, 481, 491, 570, 704]. Polynomials [36, 61, 63, 79, 114, 165, 169, 182, 185, 220, 280, 281, 292, 343, 429, 506, 515, 517, 529, 564, 579, 612, 623, 632, 666, 672, 683, 696, 718, 737, 758, 772]. Positive [35, 141, 297, 337, 378, 405, 417, 435, 615, 651]. positiver [417]. pour [222, 415, 703]. Power [231, 301, 425, 469, 472, 549, 557, 633, 672, 726, 749]. powerful [659]. powers [100, 157, 168, 273, 603, 605, 613, 770]. practical [711]. premier [468, 740]. prescribed [248, 502]. presentation [447]. primality [22]. Prime [10, 16, 34, 170, 172, 175, 176, 201, 212, 260, 285, 293, 405, 414, 418, 429, 435, 468, 506, 516, 532, 584, 647, 655, 667, 710, 721, 740, 749, 790]. primes [33, 53, 54, 121, 331, 452, 494, 604, 615, 666, 770]. Primitive [33, 114, 548]. principle [479]. Probabilistic [22]. probability [63, 564]. Problem [20, 79, 87, 213, 217, 220, 229, 396, 434, 552, 560, 568, 569, 585, 595, 629, 686, 715, 736, 770]. problème [736]. problèmes [375]. problems [62, 97, 203, 317, 375, 646, 717]. product [374, 393, 410, 486, 752]. products [392, 507]. produits [410]. Produktordnungen [374]. proendlichem [203]. progression [34, 777]. progressions [121, 622, 715, 721]. Proof [67, 158, 211, 270, 442, 723, 737]. properties [14, 29, 85, 109, 205, 307, 398, 502, 654]. property [9, 41, 418, 578, 610, 619, 695]. publications [95]. Pure [605]. Pythagoräische [42]. Pythagoras [566]. Pythagorean [42].

Quadratic [10, 17, 19, 30, 35, 40, 46, 47, 55, 66, 120, 141, 167, 174, 175, 177, 188, 196, 224, 239, 251, 255, 258, 294, 297, 303, 309, 315, 325, 331, 372, 396, 398, 404, 408, 414, 432, 445, 446, 454, 457, 459, 514, 515, 545, 566, 593, 598, 651, 656, 673, 684, 714, 724, 760, 767, 771, 773, 789, 793]. quadratische [135, 432]. quadratischen [47, 251]. quadratischer [66, 255, 325]. quartic [13, 17, 152, 299]. quaternion [64, 454, 695, 783]. quelconque [658]. Quelques [524]. question [173, 576]. questions [780]. quintic [165].

radical [113, 187]. radicals [332]. radius [682]. Ramanujan [211, 511, 634, 650]. ramification [610]. ramified [420]. Rang [98, 342]. rank [30, 98, 143, 274, 325, 342, 393, 618, 714]. ranks [747, 767]. ratio [327]. Rational [96, 136, 272, 404, 432, 444, 523, 566, 590, 591, 600, 786]. rationale [432]. rationalen [272]. rationalité [419]. Rationality [419]. rationals [13, 373, 497, 507]. Ray [372]. rayon [682]. real [9, 27, 86, 141, 251, 255, 363, 372, 396, 457, 483, 507, 684, 705, 730, 767, 773, 792]. real-quadratic [251, 255]. Realisierung [453]. realization [453, 521]. Realizations [523]. Reciprocity [11, 42, 75, 655]. recurrence [605]. recurrences [157, 481, 603, 785]. recurrent [10, 704]. récurrentes [704]. recurring [324]. recursive [689]. Rédei [56]. reduced [362, 363]. reducible [73]. Reduction [571]. réel [86]. reell [251, 255]. reell-quadratischen [251]. reell-quadratischer [255]. réels [705]. reflections [788]. regions [308]. Regular [522, 699]. Regularities [271]. regulator [345, 368]. regulators [555]. related [204, 221, 257, 284, 313, 355, 605, 725, 736, 780]. relating [345]. relation [180, 232, 689]. Relations [195, 458, 527, 537]. relationship [782]. Relative [467, 539, 624, 741]. relatives [467]. remains [222]. Remark [20, 35, 73, 125, 460]. Remarks [402, 466, 659, 664]. repeated [193, 407]. representation [51, 66, 107, 197, 221, 358, 653, 792]. representations [278, 331, 339, 446, 548, 694, 712]. Residue [398, 414, 515]. residues [10, 175, 177, 231, 303, 749]. respect [785]. restes [222]. restricted [62, 260, 374, 667]. restricting [203]. result [39, 50, 329, 376, 470]. resultant [517]. résultats [524]. results [6, 87, 433, 524, 629, 687]. Reziprozität [42]. Reziprozitätsgesetz [75]. rhymes [560]. Riemann [144, 198, 234, 237, 246, 313, 452, 528, 592, 596, 707, 728, 746]. ring [142, 268, 468, 600, 682]. rings [167, 219, 239, 249, 252, 287, 314, 391, 538, 569, 609, 619, 699]. Rogers [211, 634]. Roiter [557]. rootnumbers [64]. roots [114, 195, 213]. rotations [756, 771]. Roth [409]. rowed [59]. rules [374].

S. [156, 330, 357, 371, 511]. sa [518]. Saito [694]. Salem [286]. sans [740].

Sapiro [212]. Satz [480]. scalars [557]. Schmidt [173, 409]. Schur [21]. second [157]. Segre [616]. Selberg [250, 490, 649, 735, 738]. Self [719]. Self-similar [719]. Selmer [464]. semigroups [565]. sentence [480]. September [475, 573, 669, 764]. sequence [82, 184, 677]. sequences [14, 38, 39, 109, 161, 193, 222, 257, 271, 324, 395, 415, 491, 550, 567, 605, 660, 664, 674, 704, 706, 719]. série [419]. Series [74, 122, 147, 152, 158, 264, 301, 313, 355, 419, 426, 441, 448, 469, 472, 545, 549, 598, 618, 680, 685, 726, 757]. Serret [460]. set [377, 430, 623, 658]. Sets [31, 33, 238, 243, 276, 277, 302, 388, 436, 455, 696, 772, 779, 784]. setting [354]. seven [320]. several [79, 319]. Shidlovski [115]. Shimura [738]. short [28, 53, 532, 687]. sided [103, 218]. Siegel [217, 379, 409, 446, 618]. Sieve [15, 54, 137, 219, 428, 491]. sieves [492, 635]. Sieving [615]. sifting [54, 137]. signature [651]. similar [719]. Simple [133, 134, 246, 575]. Simultaneous [300, 331, 497, 504, 645]. simultanés [504]. singular [604]. sixth [127]. size [44]. Skolem [652]. Skubenko [8]. Small [54, 127, 137, 159, 262, 327, 513, 534, 623, 716, 724, 789]. smallest [238]. solubility [311]. solution [302, 323, 324, 652, 711, 751]. Solutions [256, 288, 375, 444]. solvability [203, 372]. Solving [571]. Some [1, 13, 14, 85, 87, 90, 97, 101, 133, 148, 164, 322, 347, 376, 394, 402, 433, 447, 452, 507, 524, 526, 538, 550, 561, 603, 605, 650, 664, 718, 779–781]. Sometimes [409]. somme [581]. Sophie [584]. space [264, 685]. special [7, 270, 271, 550]. specially [85]. specified [63]. spectrum [257]. spheres [736]. Spinor [66, 258, 548, 756, 771]. Spinorgeschlechter [66]. Spitzenformen [74]. splitting [621, 656]. splittings [55]. Square [303, 746, 754, 758]. Square-free [758]. squares [337, 601, 642, 780]. Stickelberger [368]. Stirling [535]. stochastically [488]. strange [689]. Straus [254]. Structure [48, 64, 72, 86, 340, 420, 624]. study [182, 222]. subfields [187, 299]. Subgroups [37, 344, 391]. subsequences [38]. Subsequent [32]. Subset [594]. subsets [378, 625]. substitutions [658]. successive [424]. suite [184]. suites [222, 415, 704]. sum [39, 277, 285, 321, 328, 337, 416, 448, 581, 613, 780]. sum-free [39]. summands [259, 502]. summation [423]. Summen [44, 255]. Sums [5, 25, 44, 67, 76, 100, 126, 132, 151, 158, 233, 255, 260, 295, 303, 307, 341, 355, 392, 425, 450, 486, 537, 539, 578, 594, 601, 603, 628, 647, 660, 661, 672, 710, 726, 752, 776, 791]. Sumsets [622]. surfaces [528]. Swinnerton [336]. Sylvester [50]. symbol [384, 493]. symbols [41]. symplectic [339]. Syracuse [533]. system [419]. système [419]. systems [115, 181]. Szemerédi [434].

tame [339]. tamely [420]. Tate [301]. Tauber [778]. Teilkörper [251]. term [490, 537]. terms [649, 735]. ternäre [135]. ternärer [66]. Ternary [66, 103, 210, 218, 454, 545]. testing [22, 586]. th [195]. their [182, 203, 299, 618]. theorem [8, 32, 67, 79, 84, 111, 115, 156, 182, 189, 199, 212, 225, 264, 320, 338, 359, 460, 496, 502, 557, 576, 584, 616, 643, 662, 666, 675, 721, 727, 742, 778]. theorems [83, 163, 237, 582, 583, 650, 743]. theoretic [72, 654, 680]. Theorie [74]. theory [13, 74, 94, 107, 299, 301, 310, 319, 386, 429, 452, 492, 498, 553, 592, 625, 775]. Theta [35, 147, 358, 384, 459, 545, 618]. theta-series [618]. Three [14, 96, 150, 337, 537, 601]. Three-term [537]. Thue [274, 409, 652, 711]. tiling [754]. tiny [678]. top [74]. Torsion [252, 314, 332, 538, 698, 786]. totalement [513]. totally [27, 59, 141, 513]. trace [490, 536, 649, 735]. transcendance [524, 703]. Transcendence [223, 348, 524, 534, 703]. Transcendental [110, 275, 424]. transfer [479]. Transformation [35, 77, 158, 280, 358, 541]. transformations [446]. transitive [417]. trigonometric [577]. Tripel [42]. Triples [42]. tupel [730]. Turaán [95]. Turán [93, 94, 723]. Two [14, 89, 210, 232, 277, 393, 441, 596, 654, 667, 773]. two-dimensional [654]. Two-weight [210]. type [86, 151]. types [747].

Uniform [38, 124, 376, 628]. unimodal [737]. unimodular [46, 55, 125, 618]. Union [455]. Union-free [455]. unions [641]. Uniqueness [477]. unit [98, 201, 251, 365, 366]. Unités [81, 86, 342, 705]. Units [17, 27, 81, 86, 116, 131, 143, 180, 181, 252, 287, 314, 342, 391, 398, 514, 538, 546, 562, 632, 699, 705, 713, 750]. universal [726]. universality [792]. Unramified [671]. Unsolvability [99]. Upper [31, 446]. using [255, 321, 571].

V [530]. Valeurs [122, 262]. Valuation [310]. Value [76, 237, 294, 623, 696, 761]. valued [199, 610, 611, 772]. Values [10, 28, 62, 73, 122, 255, 262, 270, 297, 407, 443, 481, 493, 528, 536, 564, 651, 661, 758]. vandiver [726, 762]. variable [596]. variables [62, 79, 319, 667]. varieties [420, 503, 520]. variety [159, 587]. various [230]. verallgemeinerten [255]. verschränkte [374]. verteilungseigenschaft [41]. Verzweigung [203]. via [367]. vierten [251]. View [310, 317]. View-obstruction [317]. Vinogradov [577]. volume [69, 117, 153, 190, 226, 265, 304, 349, 381, 411, 438, 474, 508, 542, 572, 606, 638, 668, 700, 732, 763, 794]. vom [98]. Voronoi [365, 366].

Walum [330, 357, 371, 401]. Waring [568, 569]. way [337]. wedge [312]. weight [147, 210]. weighted [629]. Weil [358]. well [14]. well-known [14]. Werte [255]. which [74, 96, 216, 377, 502]. whose [445, 718]. Wieferich [688]. Wilson [163]. within [312]. without [516, 619, 740]. Witt [354, 614, 691]. work [94]. wreath [507].

xi [234]. xi-function [234].

years [130]. yield [377].

Zahlen [47]. Zahlentwicklungen [560]. Zahlkörper [251, 255, 325]. Zahlkörpern [98]. Zerlegungsgesetze [432]. zero [308]. zero-free [308]. Zeros [144, 198, 217, 234, 246, 298, 673, 724, 789]. Zeta [40, 56, 144, 183, 237, 246, 255, 313, 319, 592, 596, 746]. zeta-function [237, 246, 592, 746]. Zetafunktionen [255]. Znám [595]. zugeordnet [74]. zur [74, 395].

References

Bhaskaran:1979:CGF

[1] M. Bhaskaran. Construction of genus field and some applications. Journal of Number Theory, 11(4):488–497, November 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/0022314X79900283.

See comment [347].

Anonymous:1980:PF

[2] Anonymous. Pages 1–139 (February 1980). Journal of Number Theory, 12(1):??, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1980:EB

[3] Anonymous. Editorial Board. Journal of Number Theory, 12(1):i, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900657.

Zassenhaus:1980:D

[4] H. J. Zassenhaus. Dedication. Journal of Number Theory, 12(1):1, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900669.

Knopp:1980:HOI

[5] Marvin I. Knopp. Hecke operators and an identity for the Dedekind sums. Journal of Number Theory, 12(1):2–9, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900670.

Balasubramanian:1980:ENR

[6] R. Balasubramanian and K. Ramachandra. Effective and noneffective results on certain arithmetical functions. Journal of Number Theory, 12 (1):10–19, February 1980. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X80900682.

Mahler:1980:SF

[7] K. Mahler. On a special function. Journal of Number Theory, 12(1):20– 26, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X80900694.

Bambah:1980:MCT

[8] R. P. Bambah and A. C. Woods. Minkowski’s conjecture for n = 5; a theorem of Skubenko. Journal of Number Theory, 12(1):27–48, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900700.

Stark:1980:PFR

[9] H. M. Stark and D. Zagier. A property of L-functions on the real line. Journal of Number Theory, 12(1):49–52, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900712.

Walum:1980:RPL

[10] Herbert Walum. A recurrent pattern in the list of quadratic residues mod a prime and in the values of the Liouville λ function. Journal of Number Theory, 12(1):53–56, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X80900724.

Moreno:1980:HRL

[11] Carlos J. Moreno. The higher reciprocity laws: an example. Journal of Number Theory, 12(1):57–70, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X80900736.

Longyear:1980:CBB

[12] Judith Q. Longyear. The circuit basis in binary matroids. Journal of Number Theory, 12(1):71–76, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X80900748.

Edgar:1980:SCT

[13] Hugh Edgar and Brian Peterson. Some contributions to the theory of cyclic quartic extensions of the rationals. Journal of Number Theory, 12(1):77–83, February 1980. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X8090075X.

Cowles:1980:SCP

[14] John Cowles. Some congruence properties of three well-known sequences: Two notes. Journal of Number Theory, 12(1):84–86, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900761.

Bressoud:1980:EPS

[15] David M. Bressoud. Extension of the partition sieve. Journal of Number Theory, 12(1):87–100, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X80900773.

Friedlander:1980:EPI

[16] John B. Friedlander. Estimates for prime ideals. Journal of Number Theory, 12(1):101–105, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X80900785.

Leonard:1980:QCC

[17] Philip A. Leonard and Kenneth S. Williams. The quartic characters of certain quadratic units. Journal of Number Theory, 12(1):106–109, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900797.

Hartung:1980:PE

[18] Paul G. Hartung. On the Pellian equation. Journal of Number Theory, 12 (1):110–112, February 1980. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X80900803.

Cowles:1980:DCN

[19] Mary J. Cowles. On the divisibility of the class number of imaginary quadratic fields. Journal of Number Theory, 12(1):113–115, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900815.

Srinivasan:1980:RGP

[20] S. Srinivasan. A remark on Goldbach’s problem. Journal of Number Theory, 12(1):116–121, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X80900827.

Morton:1980:ESM

[21] Patrick Morton. On the elgenvectors of Schur’s matrix. Journal of Number Theory, 12(1):122–127, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X80900839.

Rabin:1980:PAT

[22] Michael O. Rabin. Probabilistic algorithm for testing primality. Journal of Number Theory, 12(1):128–138, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X80900840.

Anonymous:1980:E

[23] Anonymous. Erratum. Journal of Number Theory, 12(1):139, February 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900852.

Anonymous:1980:PM

[24] Anonymous. Pages 141–282 (May 1980). Journal of Number Theory, 12 (2):??, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Sperber:1980:FAC

[25] Steven I. Sperber. On the L-functions associated with certain exponential sums. Journal of Number Theory, 12(2):141–153, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900475.

Erdos:1980:MAB

[26] Paul Erdős and Melvyn B. Nathanson. Minimal asymptotic bases for the natural numbers. Journal of Number Theory, 12(2):154–159, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900487.

Setzer:1980:UTR

[27] C. Bennett Setzer. Units over totally real C_{2}×C_{2} fields. Journal of Number Theory, 12(2):160–175, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X80900499.

Erdos:1980:VDF

[28] P. Erdős and R. R. Hall. Values of the divisor function on short intervals. Journal of Number Theory, 12(2):176–187, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900505.

Chowla:1980:CPA

[29] S. Chowla, J. Cowles, and M. Cowles. Congruence properties of Apèry numbers. Journal of Number Theory, 12(2):188–190, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900517.

Lagarias:1980:DRI

[30] J. C. Lagarias. On determining the 4-rank of the ideal class group of a quadratic field. Journal of Number Theory, 12(2):191–196, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900529.

Ullom:1980:UBD

[31] Stephen V. Ullom. Upper bounds for p-divisibility of sets of Bernoulli numbers. Journal of Number Theory, 12(2):197–200, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900530.

Lawrence:1980:STM

[32] W. R. Lawrence. Subsequent to a theorem of Markoff. Journal of Number Theory, 12(2):201–209, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X80900542.

Powell:1980:PDC

[33] Barry J. Powell. Primitive densities of certain sets of primes. Journal of Number Theory, 12(2):210–217, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X80900554.

Pomerance:1980:NLP

[34] Carl Pomerance. A note on the least prime in an arithmetic progression. Journal of Number Theory, 12(2):218–223, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900566.

Kitaoka:1980:RTF

[35] Yoshiyuki Kitaoka. A remark on the transformation formula of theta functions associated to positive definite quadratic forms. Journal of Number Theory, 12(2):224–229, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X80900578.

Rubel:1980:DPH

[36] L. A. Rubel, A. Schinzel, and H. Tverberg. On difference polynomials and hereditarily irreducible polynomials. Journal of Number Theory, 12 (2):230–235, May 1980. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X8090058X.

Dennin:1980:S

[37] Joseph B. Dennin. Subgroups of LF (2, 2^{n}). Journal of Number Theory, 12(2):236–243, May 1980. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X80900591.

Coquet:1980:UDM

[38] Jean Coquet. On the uniform distribution modulo one of subsequences of polynomial sequences II. Journal of Number Theory, 12(2):244–250, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900608.

Levine:1980:ERS

[39] Eugene Levine. An extremal result for sum-free sequences. Journal of Number Theory, 12(2):251–257, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X8090061X.

Terras:1980:MQF

[40] Audrey Terras. The minima of quadratic forms and the behavior of Epstein and Dedekind zeta functions. Journal of Number Theory, 12(2):258– 272, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X80900621.

Burde:1980:VLG

[41] Klaus Burde. Eine verteilungseigenschaft der Legendresymbole. (German) [A distribution property of the Legendre symbols]. Journal of Number Theory, 12(2):273–277, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X80900633.

Burde:1980:PTR

[42] Klaus Burde. Pythagoräische Tripel und Reziprozität in Galoisfeldern. (German) [Pythagorean triples and reciprocity in Galois fields]. Journal of Number Theory, 12(2):278–282, May 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X80900645.

Anonymous:1980:PA

[43] Anonymous. Pages 283–437 (August 1980). Journal of Number Theory, 12(3):??, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Schmidt:1980:FGS

[44] C.-G. Schmidt. Über die Führer von Gaußschen Summen als Größencharaktere. (German) [On the leaders of Gaussian sums as size characters]. Journal of Number Theory, 12(3):283–310, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900220.

Gay:1980:NBA

[45] David Gay. Normal binomials over algebraic number fields. Journal of Number Theory, 12(3):311–326, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X80900232.

Hsia:1980:CUQ

[46] J. S. Hsia. On the classification of unimodular quadratic forms. Journal of Number Theory, 12(3):327–333, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X80900244.

Petho:1980:NDA

[47] Attila Pethö. Über die nichtlineare diophantische Approximation von quadratischen, algebraischen Zahlen. (German) [On the nonlinear diophantic approximation of quadratic algebraic numbers]. Journal of Number Theory, 12(3):334–338, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X80900256.

Ewell:1980:MSO

[48] John A. Ewell. On the multiplicative structure of odd perfect numbers. Journal of Number Theory, 12(3):339–342, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900268.

Frey:1980:CIN

[49] Gerhard Frey. Cyclic isogenies and nonstandard arithmetic. Journal of Number Theory, 12(3):343–363, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8090027X.

Knight:1980:GRS

[50] M. J. Knight. A generalization of a result of Sylvester’s. Journal of Number Theory, 12(3):364–366, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X80900281.

Stevenson:1980:RIA

[51] Edie Stevenson. On the representation of integers by p-adic diagonal forms. Journal of Number Theory, 12(3):367–371, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900293.

Chowla:1980:NCC

[52] S. Chowla, J. Cowles, and M. Cowles. On the number of conjugacy classes in SL(2, Z). Journal of Number Theory, 12(3):372–377, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8090030X.

Richards:1980:NDP

[53] Ian Richards. On the normal density of primes in short intervals. Journal of Number Theory, 12(3):378–384, August 1980. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X80900311.

Erdos:1980:SSS

[54] P. Erdős and I. Z. Ruzsa. On the small sieve. I. Sifting by primes. Journal of Number Theory, 12(3):385–394, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X80900323.

Riehm:1980:JSA

[55] C. Riehm. Jordan splittings of almost unimodular integral quadratic forms. Journal of Number Theory, 12(3):395–420, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X80900335.

Kung:1980:RZF

[56] Joseph P. S. Kung, M. Ram Murty, and Gian-Carlo Rota. On the Rédei zeta function. Journal of Number Theory, 12(3):421–436, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900347.

Hensley:1980:C

[57] Douglas Hensley. Corrigendum. Journal of Number Theory, 12(3):437, August 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X80900359.

Anonymous:1980:PN

[58] Anonymous. Pages 439–558 (November 1980). Journal of Number Theory, 12(4):??, November 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Gordon:1980:MRP

[59] Basil Gordon and Lorne Houten. Multi-rowed partitions with totally distinct parts. Journal of Number Theory, 12(4):439–444, November 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900360.

Chowla:1980:NBN

[60] P. Chowla and S. Chowla. A note on Bernoulli numbers. Journal of Number Theory, 12(4):445–446, November 1980. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X80900372.

Agou:1980:FPC

[61] Simon Agou. Sur la factorisation des polynômes f(X^{p^{2r}} − aX^{p^{r}} − bX) sur un corps fini F_{p^{s}}. (French) [On the factorization of polynomials f(X^{p^{2r}} − aX^{p^{r}} − bX) on a finite field F_{p^{s}}]. Journal of Number Theory, 12(4): 447–459, November 1980. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X80900384.

Baker:1980:DPV

[62] Roger C. Baker and Wolfgang M. Schmidt. Diophantine problems in variables restricted to the values 0 and 1. Journal of Number Theory, 12 (4):460–486, November 1980. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X80900396.

Buttsworth:1980:PGP

[63] R. N. Buttsworth. On the probability that given polynomials have a specified highest common factor. Journal of Number Theory, 12(4):487–498, November 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X80900402.

Frohlich:1980:GMS

[64] A. Fröhlich. Galois module structure and rootnumbers for quaternion extensions of degree 2^{n}. Journal of Number Theory, 12(4):499–518, November 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X80900414.

Kida:1980:ECF

[65] Yûji Kida. l-Extensions of CM-fields and cyclotomic invariants. Journal of Number Theory, 12(4):519–528, November 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X80900426.

Schulze-Pillot:1980:DDS

[66] Rainer Schulze-Pillot. Darstellung durch Spinorgeschlechter ternärer quadratischer formen. (German) [Representation by spinor families of ternary quadratic forms]. Journal of Number Theory, 12(4):529–540, November 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X80900438.

Goldberg:1980:EPP

[67] Larry A. Goldberg. An elementary proof of the Petersson–Knopp theorem on Dedekind sums. Journal of Number Theory, 12(4):541–542, November 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8090044X.

Cusick:1980:EDA

[68] T. W. Cusick. Estimates for diophantine approximation constants. Journal of Number Theory, 12(4):543–556, November 1980. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X80900451.

Anonymous:1980:AIV

[69] Anonymous. Author index for volume 12. Journal of Number Theory, 12 (4):557–558, November 1980. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X80900463.

Anonymous:1981:PF

[70] Anonymous. Pages 1–121 (February 1981). Journal of Number Theory, 13(1):??, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1981:EB

[71] Anonymous. Editorial Board. Journal of Number Theory, 13(1):i, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X81900251.

Cornell:1981:GTC

[72] Gary Cornell and Michael Rosen. Group-theoretic constraints on the structure of the class group. Journal of Number Theory, 13(1):1–11, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X81900263.

Elliott:1981:RDV

[73] P. D. T. A. Elliott. A remark on the Dirichlet values of a completely reducible polynomial (modq). Journal of Number Theory, 13(1):12–17, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X81900275.

Good:1981:BTD

[74] A. Good. Beiträge zur Theorie der Dirichletreihen, die Spitzenformen zugeordnet sind. (German) [Contributions to the theory of Dirichlet series, which are assigned top forms]. Journal of Number Theory, 13(1):18–65, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X81900287.

Burde:1981:RGG

[75] Klaus Burde. Ein Reziprozitätsgesetz in Galoisfeldern. (German) [A reciprocity law in Galois fields]. Journal of Number Theory, 13(1):66–87, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X81900299.

Bressoud:1981:VGS

[76] D. M. Bressoud. On the value of Gaussian sums. Journal of Number Theory, 13(1):88–94, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X81900305.

Adler:1981:CNN

[77] Roy Adler, Michael Keane, and Meir Smorodinsky. A construction of a normal number for the continued fraction transformation. Journal of Number Theory, 13(1):95–105, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X81900317.

Valentini:1981:AHD

[78] Robert C. Valentini and Manohar L. Madan. Automorphisms and holomorphic differentials in characteristic p. Journal of Number Theory, 13 (1):106–115, February 1981. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X81900329.

Boyd:1981:KTL

[79] David W. Boyd. Kronecker’s theorem and Lehmer’s problem for polynomials in several variables. Journal of Number Theory, 13(1):116–121, February 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X81900330.

Anonymous:1981:PM

[80] Anonymous. Pages 123–270 (May 1981). Journal of Number Theory, 13 (2):??, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Mestre:1981:CEU

[81] Jean-François Mestre. Corps euclidiens, unités exceptionnelles et courbes elliptiques. (French) [Euclidean bodies, exceptional units and elliptic curves]. Journal of Number Theory, 13(2):123–137, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X81900019.

Ramshaw:1981:DSF

[82] Lyle Ramshaw. On the discrepancy of the sequence formed by the multiples of an irrational number. Journal of Number Theory, 13(2):138–175, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X81900020.

Hayashi:1981:OTI

[83] Elmer K. Hayashi. Omega theorems for the iterated additive convolution of a nonnegative arithmetic function. Journal of Number Theory, 13(2):176–191, May 1981. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X81900032.

Uhrin:1981:GMC

[84] Béla Uhrin. On a generalization of Minkowski’s convex body theorem. Journal of Number Theory, 13(2):192–209, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X81900044.

Redmond:1981:SPS

[85] Don Redmond and R. Sivaramakrishnan. Some properties of specially multiplicative functions. Journal of Number Theory, 13(2):210–227, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X81900056.

Duval:1981:SGD

[86] Dominique Duval. Sur la structure galoisienne du groupe des unités d’un corps abélien réel de type (p, p). (French) [On the Galois structure of the group of units of a real abelian body of type (p, p)]. Journal of Number Theory, 13(2):228–245, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X81900068.

Bond:1981:SRC

[87] Robert J. Bond. Some results on the capitulation problem. Journal of Number Theory, 13(2):246–254, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X8190007X.

Furno:1981:CDI

[88] Anne Ludington Furno. Cycles of differences of integers. Journal of Number Theory, 13(2):255–261, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X81900081.

Bloom:1981:IIC

[89] John R. Bloom and Frank Gerth. The Iwasawa invariant μ in the composite of two Z_{l}-extensions. Journal of Number Theory, 13(2):262–267, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X81900093.

Mahler:1981:SID

[90] K. Mahler. On some irrational decimal fractions. Journal of Number Theory, 13(2):268–269, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X8190010X.

Anonymous:1981:A

[91] Anonymous. Addendum. Journal of Number Theory, 13(2):270, May 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X81900111.

Anonymous:1981:PA

[92] Anonymous. Pages 271–419 (August 1981). Journal of Number Theory, 13(3):??, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Alpar:1981:MPT

[93] L. Alpár. In memory of Paul Turán. Journal of Number Theory, 13 (3):271–278, August 1981. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X81900123.

Pintz:1981:PTW

[94] J. Pintz. On Paul Turán’s work in number theory. Journal of Number Theory, 13(3):279–288, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X81900135.

Anonymous:1981:LPP

[95] Anonymous. List of publications of Paul Turaán. Journal of Number Theory, 13(3):289–298, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X81900147.

Parry:1981:DPW

[96] Walter R. Parry. A determination of the points which are rational over Q of three modular curves. Journal of Number Theory, 13(3):299–302, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X81900159.

Odoni:1981:CKD

[97] R. W. K. Odoni. A conjecture of Krishnamurthy on decimal periods and some allied problems. Journal of Number Theory, 13(3):303–319, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X81900160.

Brunotte:1981:MKE

[98] Horst Brunotte and Franz Halter-Koch. Metrische Kennzeichnung von Erzeugenden für Einheitengruppen vom Rang 1 oder 2 in algebraischen Zahlkörpern. (German) [Metric marking of generators for unit groups of rank 1 or 2 in algebraic number fields]. Journal of Number Theory, 13 (3):320–333, August 1981. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X81900172.

Kliorys:1981:UBF

[99] Constantine Kliorys. Unsolvability of binary forms. Journal of Number Theory, 13(3):334–336, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X81900184.

Bremner:1981:GAE

[100] Andrew Bremner. A geometric approach to equal sums of fifth powers. Journal of Number Theory, 13(3):337–354, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X81900196.

Leshchiner:1981:SNI

[101] D. Leshchiner. Some new identities for ζ(k). Journal of Number Theory, 13(3):355–362, August 1981. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X81900202.

Galovich:1981:CNC

[102] Steven Galovich and Michael Rosen. The class number of cyclotomic function fields. Journal of Number Theory, 13(3):363–375, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X81900214.

Jackson:1981:OSI

[103] T. H. Jackson. One-sided inequalities for ternary forms, I. Journal of Number Theory, 13(3):376–397, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X81900226.

Thomas:1981:DEA

[104] E. Thomas and A. T. Vasquez. Diophantine equations arising from cubic number fields. Journal of Number Theory, 13(3):398–414, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X81900238.

Mimura:1981:OGI

[105] Yoshio Mimura. On orthogonal groups over the integers of a non-dyadic local field. Journal of Number Theory, 13(3):415–419, August 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8190024X.

Anonymous:1981:PN

[106] Anonymous. Pages 421–563 (November 1981). Journal of Number Theory, 13(4):??, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Gurak:1981:RTF

[107] S. Gurak. On the representation theory for full decomposable forms. Journal of Number Theory, 13(4):421–442, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X81900342.

Ayoub:1981:EP

[108] R. G. Ayoub and S. Chowla. On Euler’s polynomial. Journal of Number Theory, 13(4):443–445, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X81900354.

Burr:1981:CPP

[109] Stefan A. Burr and Paul Erdős. Completeness properties of perturbed sequences. Journal of Number Theory, 13(4):446–455, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X81900366.

Nettler:1981:TCF

[110] Gideon Nettler. Transcendental continued fractions. Journal of Number Theory, 13(4):456–462, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X81900378.

Hinz:1981:TBD

[111] Jürgen G. Hinz. On the theorem of Barban and Davenport–Halberstam in algebraic number fields. Journal of Number Theory, 13(4):463–484, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X8190038X.

Delang:1981:E

[112] Li Delang. On the equation 4n = 1x + 1y + 1z. Journal of Number Theory, 13(4):485–494, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X81900391.

Newman:1981:RDE

[113] Morris Newman. A radical diophantine equation. Journal of Number Theory, 13(4):495–498, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X81900408.

Madden:1981:PPR

[114] Daniel J. Madden. Polynomials and primitive roots in finite fields. Journal of Number Theory, 13(4):499–514, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8190041X.

Osgood:1981:NPS

[115] Charles F. Osgood. Nearly perfect systems and effective generalizations of Shidlovski’s Theorem. Journal of Number Theory, 13(4):515–540, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X81900421.

Kleinert:1981:EGU

[116] Ernst Kleinert. Einheiten in Z[D_{2m}]. (German) [Units in Z[D_{2m}]]. Journal of Number Theory, 13(4):541–561, November 1981. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X81900433.

Anonymous:1981:AIV

[117] Anonymous. Author index for volume 13. Journal of Number Theory, 13 (4):562–563, November 1981. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X81900445.

Anonymous:1982:PF

[118] Anonymous. Pages 1–120 (February 1982). Journal of Number Theory, 14(1):??, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1982:EBa

[119] Anonymous. Editorial Board. Journal of Number Theory, 14(1):i, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900531.

Earnest:1982:PGQ

[120] A. G. Earnest. Partitionings of a genus of quadratic forms. Journal of Number Theory, 14(1):1–8, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X82900543.

Grosswald:1982:APC

[121] Emil Grosswald. Arithmetic progressions that consist only of primes. Journal of Number Theory, 14(1):9–31, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900555.

Cassou-Nogues:1982:VEN

[122] Pierrette Cassou-Noguès. Valeurs aux entiers négatifs des séries de Dirichlet associées a un polynôme, I. (French) [Values with negative integers in the Dirichlet series associated with a polynomial, I]. Journal of Number Theory, 14(1):32–64, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X82900567.

Chowla:1982:BNI

[123] P. Chowla and S. Chowla. On Bernoulli numbers, II. Journal of Number Theory, 14(1):65–66, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X82900579.

Baumert:1982:UC

[124] L. D. Baumert, W. H. Mills, and Robert L. Ward. Uniform cyclotomy. Journal of Number Theory, 14(1):67–82, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900580.

Schulze-Pillot:1982:RNH

[125] Rainer Schulze-Pillot. Remark on a note of Hsia about classification of unimodular lattices. Journal of Number Theory, 14(1):83–85, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900592.

Alladi:1982:AES

[126] Krishnaswami Alladi. Asymptotic estimates of sums involving the Moebius function. Journal of Number Theory, 14(1):86–98, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900609.

Pohst:1982:CNF

[127] Michael Pohst. On the computation of number fields of small discriminants including the minimum discriminants of sixth degree fields. Journal of Number Theory, 14(1):99–117, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900610.

Leopoldt:1982:O

[128] H. W. Leopoldt. Obituary. Journal of Number Theory, 14(1):118–120, February 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X82900622.

Anonymous:1982:PAa

[129] Anonymous. Pages 121–279 (April 1982). Journal of Number Theory, 14 (2):??, April 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Mahler:1982:FYM

[130] K. Mahler. Fifty years as a mathematician. Journal of Number Theory, 14(2):121–155, April 1982. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X82900440.

Galovich:1982:UCG

[131] Steven Galovich and Michael Rosen. Units and class groups in cyclotomic function fields. Journal of Number Theory, 14(2):156–184, April 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900452.

MacLeod:1982:FPS

[132] R. A. MacLeod. Fractional part sums and divisor functions. Journal of Number Theory, 14(2):185–227, April 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900464.

Shallit:1982:SCF

[133] J. O. Shallit. Simple continued fractions for some irrational numbers, II. Journal of Number Theory, 14(2):228–231, April 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900476.

Petho:1982:SCF

[134] Attila Pethö. Simple continued fractions for the Fredholm numbers. Journal of Number Theory, 14(2):232–236, April 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900488.

Schulze-Pillot:1982:DDD

[135] Rainer Schulze-Pillot. Darstellung durch definite ternäre quadratische Formen. Journal of Number Theory, 14(2):237–250, April 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8290049X.

Gurak:1982:REF

[136] S. Gurak. On the rational equivalence of full decomposable forms. Journal of Number Theory, 14(2):251–259, April 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900506.

Ruzsa:1982:SSI

[137] Imre Z. Ruzsa. On the small sieve. II. sifting by composite numbers. Journal of Number Theory, 14(2):260–268, April 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900518.

Ramharter:1982:ADA

[138] Gerhard Ramharter. Über asymmetrische diophantische Approximationen. (German) [On asymmetric diophantine approximations]. Journal of Number Theory, 14(2):269–279, April 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8290052X.

Anonymous:1982:PJ

[139] Anonymous. Pages 281–424 (June 1982). Journal of Number Theory, 14 (3):??, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Frank:1982:OP

[140] Evelyn Frank. Oskar Perron (1880–1975). Journal of Number Theory, 14 (3):ii, 281–291, June 1982. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X82900634.

Dress:1982:ITP

[141] Andreas Dress and Rudolf Scharlau. Indecomposable totally positive numbers in real quadratic orders. Journal of Number Theory, 14(3):292–306, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900646.

Guralnick:1982:IMU

[142] Robert M. Guralnick. Isomorphism of modules under ground ring extensions. Journal of Number Theory, 14(3):307–314, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900658.

Feng:1982:RGC

[143] Keqin Feng. The rank of group of cyclotomic units in abelian fields. Journal of Number Theory, 14(3):315–326, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8290066X.

Mueller:1982:DBC

[144] Julia Mueller. On the difference between consecutive zeros of the Riemann zeta function. Journal of Number Theory, 14(3):327–331, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900671.

Koblitz:1982:CBN

[145] Neal Koblitz. On Carlitz’s q-Bernoulli numbers. Journal of Number Theory, 14(3):332–339, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X82900683.

Kida:1982:CEF

[146] Yûji Kida. Cyclotomic Z_{2}-extensions of J-fields. Journal of Number Theory, 14(3):340–352, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X82900695.

Sturm:1982:TSW

[147] Jacob Sturm. Theta series of weight 32. Journal of Number Theory, 14(3):353–361, June 1982. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X82900701.

Gessel:1982:SCA

[148] Ira Gessel. Some congruences for Apéry numbers. Journal of Number Theory, 14(3):362–368, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X82900713.

Chowla:1982:X

[149] S. Chowla, J. Cowles, and M. Cowles. On x^{3}+y^{3} = D. Journal of Number Theory, 14(3):369–373, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X82900725.

Stevenson:1982:ACT

[150] Edie Stevenson. The Artin conjecture for three diagonal cubic forms. Journal of Number Theory, 14(3):374–390, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900737.

Apostol:1982:ISD

[151] Tom M. Apostol and Thiennu H. Vu. Identities for sums of Dedekind type. Journal of Number Theory, 14(3):391–396, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900749.

Shanks:1982:DQA

[152] Daniel Shanks. Dihedral quartic approximations and series for π. Journal of Number Theory, 14(3):397–423, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900750.

Anonymous:1982:AIVa

[153] Anonymous. Author index for volume 14. Journal of Number Theory, 14 (3):424, June 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X82900762.

Anonymous:1982:PAb

[154] Anonymous. Pages 1–142 (August 1982). Journal of Number Theory, 15 (1):??, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1982:EBb

[155] Anonymous. Editorial Board. Journal of Number Theory, 15(1):i, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900774.

Wang:1982:TC

[156] Kai Wang. On a theorem of S. Chowla. Journal of Number Theory, 15(1): 1–4, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X82900786.

Petho:1982:PPS

[157] Attila Pethö. Perfect powers in second order linear recurrences. Journal of Number Theory, 15(1):5–13, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X82900798.

Apostol:1982:EPT

[158] Tom M. Apostol. Elementary proof of the transformation formula for Lambert series involving generalized Dedekind sums. Journal of Number Theory, 15(1):14–24, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X82900804.

Shimura:1982:MAV

[159] Goro Shimura. Models of an abelian variety with complex multiplication over small fields. Journal of Number Theory, 15(1):25–35, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900816.

Hafner:1982:AOC

[160] James Lee Hafner. On the average order of a class of arithmetical functions. Journal of Number Theory, 15(1):36–76, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900828.

Perelli:1982:PS

[161] A. Perelli and U. Zannier. On periodic mod p sequences. Journal of Number Theory, 15(1):77–82, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X8290083X.

Conway:1982:ELD

[162] J. H. Conway and N. J. A. Sloane. On the enumeration of lattices of determinant one. Journal of Number Theory, 15(1):83–94, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900841.

Chao:1982:GTW

[163] Chong-Yun Chao. Generalizations of theorems of Wilson, Fermat and Euler. Journal of Number Theory, 15(1):95–114, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900853.

Erdos:1982:SAF

[164] P. Erdős and A. Sárközy. Some asymptotic formulas on generalized divisor functions, II. Journal of Number Theory, 15(1):115–136, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900865.

Roland:1982:PFQ

[165] G. Roland, N. Yui, and D. Zagier. A parametric family of quintic polynomials with Galois group D_{5}. Journal of Number Theory, 15(1):137–142, August 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X82900877.

Anonymous:1982:PO

[166] Anonymous. Pages 143–282 (October 1982). Journal of Number Theory, 15(2):??, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Lazard:1982:MAR

[167] Daniel Lazard. On the minimal algorithm in rings of imaginary quadratic integers. Journal of Number Theory, 15(2):143–148, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8290021X.

Tsang:1982:DIM

[168] Kai-Man Tsang. Diophantine inequalities with mixed powers. Journal of Number Theory, 15(2):149–163, October 1982. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900221.

Gyory:1982:ICP

[169] K. Györy. On the irreducibility of a class of polynomials, III. Journal of Number Theory, 15(2):164–181, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900233.

Williams:1982:CMC

[170] Kenneth S. Williams. Congruences modulo 8 for the class numbers of Q(\sqrt{±p}), p ≡ 3(mod4) a prime. Journal of Number Theory, 15(2):182– 198, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X82900245.

Kenku:1982:NIC

[171] M. A. Kenku. On the number of Q-isomorphism classes of elliptic curves in each Q-isogeny class. Journal of Number Theory, 15(2):199–202, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900257.

Pleasants:1982:NPF

[172] P. A. B. Pleasants. The number of prime factors of binomial coefficients. Journal of Number Theory, 15(2):203–225, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900269.

Freiling:1982:AQS

[173] Chris Freiling. An answer to a question of Schmidt on (α, β) games. Journal of Number Theory, 15(2):226–228, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900270.

Sarnak:1982:CNI

[174] Peter Sarnak. Class numbers of indefinite binary quadratic forms. Journal of Number Theory, 15(2):229–247, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900282.

Walum:1982:DQR

[175] Herbert Walum. On the distribution of quadratic residues modulo a prime. Journal of Number Theory, 15(2):248–251, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900294.

Bentz:1982:DDP

[176] Hans J. Bentz. Discrepancies in the distribution of prime numbers. Journal of Number Theory, 15(2):252–274, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900300.

Sarges:1982:LQN

[177] Heidrun Sarges and Werner Schaal. Least quadratic non-residues in algebraic number fields. Journal of Number Theory, 15(2):275–281, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900312.

Delang:1982:LE

[178] Li Delang. Letter to the editor. Journal of Number Theory, 15(2):282, October 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X82900324.

Anonymous:1982:PD

[179] Anonymous. Pages 283–424 (December 1982). Journal of Number Theory, 15(3):??, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Appelgate:1982:PEM

[180] H. Appelgate and H. Onishi. Periodic expansion of modules and its relation to units. Journal of Number Theory, 15(3):283–294, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900336.

Frei:1982:FSU

[181] Günther Frei. Fundamental systems of units in biquadratic parametric number fields. Journal of Number Theory, 15(3):295–303, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X82900348.

Thaine:1982:PGB

[182] F. Thaine. Polynomials generalizing binomial coefficients and their application to the study of Fermat’s last theorem. Journal of Number Theory, 15(3):304–317, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X8290035X.

Kato:1982:ZME

[183] Goro C. Kato and Saul Lubkin. Zeta matrices of elliptic curves. Journal of Number Theory, 15(3):318–330, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900361.

Tenenbaum:1982:DDS

[184] Gérald Tenenbaum. Sur la densité divisorielle d’une suite d’entiers. (French) [On the density of a sequence of integers]. Journal of Number Theory, 15(3):331–346, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X82900373.

Jensen:1982:PGG

[185] Christian U. Jensen and Noriko Yui. Polynomials with D_{p} as Galois group. Journal of Number Theory, 15(3):347–375, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900385.

Tzanakis:1982:DE

[186] Nicholas Tzanakis. The diophantine equation x^{3} + 3y^{3} = 2^{n}. Journal of Number Theory, 15(3):376–387, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X82900397.

deOrozco:1982:LSR

[187] MariáAcosta de Orozco and William Yslas Vélez. The lattice of subfields of a radical extension. Journal of Number Theory, 15(3):388–405, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X82900403.

Kneser:1982:CBQ

[188] Martin Kneser. Composition of binary quadratic forms. Journal of Number Theory, 15(3):406–413, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X82900415.

Agoh:1982:FLT

[189] Takashi Agoh. On Fermat’s last theorem and the Bernoulli numbers. Journal of Number Theory, 15(3):414–422, December 1982. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X82900427.

Anonymous:1982:AIVb

[190] Anonymous. Author index for volume 15. Journal of Number Theory, 15 (3):423–424, December 1982. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X82900439.

Anonymous:1983:PF

[191] Anonymous. Pages 1–146 (February 1983). Journal of Number Theory, 16(1):??, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1983:EBa

[192] Anonymous. Editorial Board. Journal of Number Theory, 16(1):i, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900252.

Slater:1983:RDS

[193] Peter J. Slater. On repeated difference sequences of a permutation of N. Journal of Number Theory, 16(1):1–5, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900264.

Nart:1983:EAG

[194] Enric Nart and Núria Vila. Equations with absolute Galois group isomorphic to A_{n}. Journal of Number Theory, 16(1):6–13, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900276.

Richman:1983:PRA

[195] David R. Richman. Polynomial relations among the n-th roots of one. Journal of Number Theory, 16(1):14–18, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900288.

Nicholls:1983:MIB

[196] Peter J. Nicholls. The minima of indefinite binary quadratic forms. Journal of Number Theory, 16(1):19–30, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8390029X.

Lewittes:1983:CDR

[197] Joseph Lewittes. Characters and decomposition of a representation in a number field. Journal of Number Theory, 16(1):31–48, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900306.

Conrey:1983:ZDRa

[198] Brian Conrey. Zeros of derivatives of Riemann’s ξ-function on the critical line. Journal of Number Theory, 16(1):49–74, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900318.

Stawski:1983:THN

[199] M. S. Stawski. On the theorem of Hartogs in non–Archimedean valued fields. Journal of Number Theory, 16(1):75–86, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8390032X.

Porubsky:1983:FCI

[200] Štefan Porubský. Further congruences involving Bernoulli numbers. Journal of Number Theory, 16(1):87–94, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900331.

Brown:1983:CNF

[201] Ezra Brown. The class number and fundamental unit of Q(\sqrt{2p}), for p ≡ 1(mod16) a prime. Journal of Number Theory, 16(1):95–99, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900343.

Hudson:1983:LE

[202] Richard H. Hudson and Kenneth S. Williams. On Legendre’s equation ax^{2} + by^{2} + cz^{2} = 0. Journal of Number Theory, 16(1):100–105, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900355.

Klingen:1983:EPK

[203] Norbert Klingen. Einbettungsprobleme mit proendlichem Kern und ihre Lösbarkeit bei Beschränkung der Verzweigung. (German) [Embedding problems with a finite core and their solvability by restricting the branching]. Journal of Number Theory, 16(1):106–118, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900367.

Ivic:1983:NAG

[204] A. Ivić. On the number of Abelian groups of a given order and on certain related multiplicative functions. Journal of Number Theory, 16(1): 119–137, February 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900379.

Mimura:1983:CPA

[205] Yoshio Mimura. Congruence properties of Apéry numbers. Journal of Number Theory, 16(1):138–146, February 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900380.

Anonymous:1983:PAa

[206] Anonymous. Pages 147–284 (April 1983). Journal of Number Theory, 16 (2):??, April 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Murty:1983:AC

[207] M. Ram Murty. On Artin’s conjecture. Journal of Number Theory, 16 (2):147–168, April 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900392.

Schmidt:1983:DAE

[208] Asmus L. Schmidt. Diophantine approximation in the Eisensteinian field. Journal of Number Theory, 16(2):169–204, April 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900409.

Wolfskill:1983:EBF

[209] J. Wolfskill. On equivalence of binary forms. Journal of Number Theory, 16(2):205–211, April 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900410.

Bremner:1983:TWT

[210] A. Bremner, R. Calderbank, P. Hanlon, P. Morton, and J. Wolfskill. Twoweight ternary codes and the equation y^{2} = 4×3^{a}+13. Journal of Number Theory, 16(2):212–234, April 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X83900422.

Bressoud:1983:EPR

[211] D. M. Bressoud. An easy proof of the Rogers–Ramanujan identities. Journal of Number Theory, 16(2):235–241, April 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900434.

Heath-Brown:1983:PSP

[212] D. R. Heath-Brown. The Pjateckiĭ–Šapiro prime number theorem. Journal of Number Theory, 16(2):242–266, April 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900446.

Elliott:1983:DRC

[213] P. D. T. A. Elliott. On the distribution of the roots of certain congruences and a problem for additive arithmetic functions. Journal of Number Theory, 16(2):267–282, April 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X83900458.

Buttsworth:1983:C

[214] Robert Buttsworth and Peter Sarnak. Corrigenda. Journal of Number Theory, 16(2):283–284, April 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X8390046X.

Anonymous:1983:PJ

[215] Anonymous. Pages 285–431 (June 1983). Journal of Number Theory, 16 (3):??, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Elliott:1983:AAF

[216] P. D. T. A. Elliott. On additive arithmetic functions f(n) for which f(an + b) − f(cn + d) is bounded. Journal of Number Theory, 16(3):285– 310, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X83900598.

Balasubramanian:1983:SZG

[217] R. Balasubramanian and C. J. Mozzochi. Siegel zeros and the Goldbach problem. Journal of Number Theory, 16(3):311–332, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900604.

Jackson:1983:OSI

[218] T. H. Jackson. One-sided inequalities for ternary forms: II. Journal of Number Theory, 16(3):333–342, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X83900616.

Webb:1983:SMP

[219] William A. Webb. Sieve methods for polynomial rings over finite fields. Journal of Number Theory, 16(3):343–355, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900628.

Lawton:1983:PBC

[220] Wayne M. Lawton. A problem of Boyd concerning geometric means of polynomials. Journal of Number Theory, 16(3):356–362, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8390063X.

Coquet:1983:GCR

[221] Jean Coquet and Hubert Delange. Graphes connexes représentation des entiers et équirépartition. (French) [Related graph representation of integers and equirepartition]. Journal of Number Theory, 16(3):363–375, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900641.

Faure:1983:ERP

[222] Henri Faure. Étude des restes pour les suites de Van Der Corput généralisées. (French) [Study of the remains of generalized Van Der Corput sequences]. Journal of Number Theory, 16(3):376–394, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900653.

Escassut:1983:TO

[223] Alain Escassut. Transcendence order over Q_{p} in C_{p}. Journal of Number Theory, 16(3):395–402, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X83900665. See comment [348].

Bambah:1983:CJN

[224] R. P. Bambah, V. C. Dumir, and R. J. Hans-Gill. On a conjecture of Jackson on non-homogeneous quadratic forms. Journal of Number Theory, 16(3):403–419, June 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900677.

Mollin:1983:CNG

[225] R. A. Mollin. Class numbers and a generalized Fermat theorem. Journal of Number Theory, 16(3):420–429, June 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X83900689.

Anonymous:1983:AIVa

[226] Anonymous. Author index for volume 16. Journal of Number Theory, 16(3):430–431, June 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900690.

Anonymous:1983:PAb

[227] Anonymous. Pages 1–133 (August 1983). Journal of Number Theory, 17 (1):??, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1983:EBb

[228] Anonymous. Editorial Board. Journal of Number Theory, 17(1):i, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8390001X.

Canfield:1983:POC

[229] E. R. Canfield, Paul Erdős, and Carl Pomerance. On a problem of Oppenheim concerning “factorisatio numerorum”. Journal of Number Theory, 17(1):1–28, August 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900021.

Murty:1983:DVC

[230] M. Ram Murty and V. Kumar Murty. On the density of various classes of groups. Journal of Number Theory, 17(1):29–36, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900033.

Lehmer:1983:PPR

[231] Emma Lehmer. Patterns of power residues. Journal of Number Theory, 17(1):37–46, August 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900045.

Ryde:1983:RBT

[232] Folke Ryde. On the relation between two Minkowski functions. Journal of Number Theory, 17(1):47–51, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X83900057.

Hinz:1983:CSA

[233] Jürgen G. Hinz. Character sums in algebraic number fields. Journal of Number Theory, 17(1):52–70, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X83900069.

Conrey:1983:ZDRb

[234] Brian Conrey. Zeros of derivatives of Riemann’s xi-function of the critical line. II. Journal of Number Theory, 17(1):71–75, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900070.

Turk:1983:NFE

[235] Jan Turk. A note on the Fermat equation. Journal of Number Theory, 17(1):76–79, August 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900082.

Cook:1983:PAE

[236] R. J. Cook. Pairs of additive equations II. Large odd degree. Journal of Number Theory, 17(1):80–92, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X83900094.

Ghosh:1983:RZF

[237] A. Ghosh. On the Riemann zeta-function — Mean value theorems and the distribution of |S(T)|. Journal of Number Theory, 17(1):93–102, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900100.

Nathanson:1983:LSM

[238] Melvyn B. Nathanson. Largest and smallest maximal sets of pairwise disjoint partitions. Journal of Number Theory, 17(1):103–112, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900112.

Parimala:1983:QFP

[239] Raman Parimala. Quadratic forms over polynomial rings over global fields. Journal of Number Theory, 17(1):113–115, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900124.

Feng:1983:AHM

[240] Xuning Feng. An analog of η(z) in the Hilbert modular case. Journal of Number Theory, 17(1):116–126, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900136.

Shorey:1983:DCC

[241] T. N. Shorey. Divisors of convergents of a continued fraction. Journal of Number Theory, 17(1):127–133, August 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900148.

Anonymous:1983:PO

[242] Anonymous. Pages 135–283 (October 1983). Journal of Number Theory, 17(2):??, October 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Abbott:1983:ACS

[243] H. L. Abbott. On the additive completion of sets of integers. Journal of Number Theory, 17(2):135–143, October 1983. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8390015X.

Tzanakis:1983:DE

[244] Nicholas Tzanakis. On the diophantine equation y^{2} − D = 2^{k}. Journal of Number Theory, 17(2):144–164, October 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900161.

Mollin:1983:CP

[245] R. A. Mollin. On the cyclotomic polynomial. Journal of Number Theory, 17(2):165–175, October 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900173.

Anderson:1983:SZR

[246] Robert J. Anderson. Simple zeros of the Riemann zeta-function. Journal of Number Theory, 17(2):176–182, October 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900185.

Hafner:1983:DAO

[247] James L. Hafner. The distribution and average order of the coefficients of Dedekind ζ functions. Journal of Number Theory, 17(2):183–190, October 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900197.

Gerth:1983:ABN

[248] Frank Gerth. Asymptotic behavior of number fields with prescribed ℓ-class numbers. Journal of Number Theory, 17(2):191–203, October 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900203.

Ayoub:1983:CBI

[249] Christine W. Ayoub. On constructing bases for ideals in polynomial rings over the integers. Journal of Number Theory, 17(2):204–225, October 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900215.

Moreno:1983:CSF

[250] Carlos Julio Moreno. The Chowla–Selberg formula. Journal of Number Theory, 17(2):226–245, October 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X83900227.

Stender:1983:EPZ

[251] Hans-Joachim Stender. Einheitenbasen für parametrisierte Zahlkörper vierten und achten Grades mit beliebigem reell-quadratischen Teilkörper. (German) [Unit bases for parameterized number fields of fourth and eighth degree with any real-quadratic partial body]. Journal of Number Theory, 17(2):246–269, October 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900239.

Luthar:1983:TUI

[252] I. S. Luthar and A. K. Bhandari. Torsion units of integral group rings of metacyclic groups. Journal of Number Theory, 17(2):270–283, October 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900240.

Anonymous:1983:PD

[253] Anonymous. Pages 285–418 (December 1983). Journal of Number Theory, 17(3):??, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1983:EGS

[254] Anonymous. Ernst G. Straus: February 25, 1922–July 12, 1983. Journal of Number Theory, 17(3):iv–v, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X83900471.

Halsritter:1983:BWV

[255] Ulrich Halsritter. Berechnung der Werte von verallgemeinerten Zetafunktionen reell-quadratischer Zahlkörper mittels dedekindscher summen. (German) [Calculation of the values of generalized zeta functions of realquadratic number fields using Dedekind sums]. Journal of Number Theory, 17(3):285–322, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X83900483.

Mohanty:1983:NCI

[256] S. P. Mohanty and A. M. S. Ramasamy. On the number of coprime integral solutions of y ^{2} = x ^{3} + k. Journal of Number Theory, 17(3):323–326, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X83900495.

Coquet:1983:FBS

[257] Jean Coquet, Georges Rhin, and Philippe Toffin. Fourier–Bohr spectrum of sequences related to continued fractions. Journal of Number Theory, 17 (3):327–336, December 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900501.

Benham:1983:SEQ

[258] J. W. Benham and J. S. Hsia. Spinor equivalence of quadratic forms. Journal of Number Theory, 17(3):337–342, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900513.

Sonn:1983:DSC

[259] Jack Sonn. Direct summands of class groups. Journal of Number Theory, 17(3):343–349, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X83900525.

Hazlewood:1983:SNR

[260] Donald G. Hazlewood. Sums over numbers with restricted prime factors. Journal of Number Theory, 17(3):350–365, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X83900537.

Drobot:1983:ACP

[261] V. Drobot. Approximation of curves by polygons. Journal of Number Theory, 17(3):366–374, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X83900549.

Nicolas:1983:PVF

[262] Jean-Louis Nicolas. Petites valeurs de la fonction d’Euler. (French) [Small values of the Euler function]. Journal of Number Theory, 17(3):375–388, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X83900550.

Rumely:1983:FGP

[263] Robert S. Rumely. A formula for the grössencharacter of a parametrized elliptic curve. Journal of Number Theory, 17(3):389–402, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900562.

Mahler:1983:TGN

[264] Kurt Mahler. On a theorem in the geometry of numbers in a space of Laurent series. Journal of Number Theory, 17(3):403–416, December 1983. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X83900574.

Anonymous:1983:AIVb

[265] Anonymous. Author index for volume 17. Journal of Number Theory, 17 (3):417–418, December 1983. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X83900586.

Anonymous:1984:PF

[266] Anonymous. Pages 1–134 (February 1984). Journal of Number Theory, 18(1):??, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1984:EBa

[267] Anonymous. Editorial Board. Journal of Number Theory, 18(1):i, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900374.

Cantor:1984:DER

[268] David C. Cantor and Peter Roquette. On Diophantine equations over the ring of all algebraic integers. Journal of Number Theory, 18(1):1–26, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X84900386.

Ruzsa:1984:GMA

[269] Imre Z. Ruzsa. Generalized moments of additive functions. Journal of Number Theory, 18(1):27–33, February 1984. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84900398.

Powell:1984:PIF

[270] Barry J. Powell. Proof of the impossibility of the Fermat equation X^{p} + Y^{p} = Z^{p} for special values of p and of the more general equation bX^{n} + cY^{n} = dZ^{n}. Journal of Number Theory, 18(1):34–40, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900404.

Hellekalek:1984:RDS

[271] Peter Hellekalek. Regularities in the distribution of special sequences. Journal of Number Theory, 18(1):41–55, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900416.

Warkentin:1984:NRF

[272] Peter Warkentin. Die Normformgleichung über dem rationalen Funktionenkörper. (German) [The norm form equation over the rational function field]. Journal of Number Theory, 18(1):56–68, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900428.

Baker:1984:DIM

[273] R. C. Baker and G. Harman. Diophantine inequalities with mixed powers. Journal of Number Theory, 18(1):69–85, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8490043X.

Chowla:1984:TRC

[274] S. Chowla, J. Cowles, and M. Cowles. On Y ^{2} = X ^{3} + k and the Thue rank of cubic curves. Journal of Number Theory, 18(1):86–90, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900441.

Bundschuh:1984:TCF

[275] Peter Bundschuh. Transcendental continued fractions. Journal of Number Theory, 18(1):91–98, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X84900453.

Erdos:1984:DSD

[276] P. Erdős and Robert Freud. On disjoint sets of differences. Journal of Number Theory, 18(1):99–109, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84900465.

Olson:1984:STS

[277] John E. Olson. On the sum of two sets in a group. Journal of Number Theory, 18(1):110–120, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X84900477.

Kirschenhofer:1984:DDC

[278] P. Kirschenhofer and R. F. Tichy. On the distribution of digits in Cantor representations of integers. Journal of Number Theory, 18(1):121–134, February 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X84900489.

Anonymous:1984:PAa

[279] Anonymous. Pages 135–239 (April 1984). Journal of Number Theory, 18 (2):??, April 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Jutila:1984:TFD

[280] Matti Jutila. Transformation formulae for Dirichlet polynomials. Journal of Number Theory, 18(2):135–156, April 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84900490.

Heidrich:1984:FEP

[281] Arnošt J. J. Heidrich. On the factorization of Eulerian polynomials. Journal of Number Theory, 18(2):157–168, April 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84900507.

Guralnick:1984:GM

[282] Robert M. Guralnick. The genus of a module. Journal of Number Theory, 18(2):169–177, April 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900519.

Murty:1984:NGG

[283] M. Ram Murty and V. Kumar Murty. On the number of groups of a given order. Journal of Number Theory, 18(2):178–191, April 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900520.

Tzanakis:1984:DER

[284] Nicholas Tzanakis. The diophantine equation x^{3} − 3xy^{2} − y^{3} = 1 and related equations. Journal of Number Theory, 18(2):192–205, April 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900532.

Hensley:1984:SIA

[285] Douglas Hensley. The sum of a^{Ω(n)} over integers n ≤ x with all prime factors between a and y. Journal of Number Theory, 18(2):206–212, April 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900544.

Chinburg:1984:SNF

[286] T. Chinburg. Salem numbers and L-functions. Journal of Number Theory, 18(2):213–214, April 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900556.

Bhandari:1984:CCC

[287] Ashwani K. Bhandari and Indar S. Luthar. Certain conjugacy classes of units in integral group rings of metacyclic groups. Journal of Number Theory, 18(2):215–228, April 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X84900568.

Ng:1984:NS

[288] Eugene Kwan-Sang Ng. On the number of solutions of N − p = P_{3}. Journal of Number Theory, 18(2):229–237, April 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8490057X.

Mollin:1984:C

[289] Richard Mollin. Corrigenda. Journal of Number Theory, 18(2):238–239, April 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900581.

Anonymous:1984:PJ

[290] Anonymous. Pages 241–397 (June 1984). Journal of Number Theory, 18 (3):??, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Murty:1984:AAC

[291] M. Ram Murty. An analogue of Artin’s conjecture for Abelian extensions. Journal of Number Theory, 18(3):241–248, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900593.

Matthews:1984:PPA

[292] R. Matthews. Permutation polynomials over algebraic number fields. Journal of Number Theory, 18(3):249–260, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8490060X.

Cohen:1984:PNF

[293] Daniel I. A. Cohen and Talbot M. Katz. Prime numbers and the first digit phenomenon. Journal of Number Theory, 18(3):261–268, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900611.

Moreno:1984:VAF

[294] Carlos J. Moreno. The value of L(12, χ) for Abelian L-functions of complex quadratic fields. Journal of Number Theory, 18(3):269–288, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900623.

Dieter:1984:CSF

[295] Ulrich Dieter. Cotangent sums, a further generalization of Dedekind sums. Journal of Number Theory, 18(3):289–305, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900635.

Wang:1984:MD

[296] Kai Wang. On Maillet determinant. Journal of Number Theory, 18 (3):306–312, June 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900647.

Bambah:1984:PVN

[297] R. P. Bambah, V. C. Dumir, and R. J. Hans-Gill. Positive values of nonhomogeneous indefinite quadratic forms II. Journal of Number Theory, 18(3):313–341, June 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900659.

Brownawell:1984:AZF

[298] W. Dale Brownawell. On p-adic zeros of forms. Journal of Number Theory, 18(3):342–349, June 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900660.

Xianke:1984:CQF

[299] Zhang Xianke. Cyclic quartic fields and genus theory of their subfields. Journal of Number Theory, 18(3):350–355, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900672.

Mohanty:1984:SDE

[300] S. P. Mohanty and A. M. S. Ramasamy. The simultaneous diophantine equations 5Y^{2} − 20 = X^{2} and 2Y^{2} + 1 = Z^{2}. Journal of Number Theory, 18(3):356–359, June 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900684.

Sekiguchi:1984:LTT

[301] Koji Sekiguchi. The Lubin–Tate theory for formal power series fields with finite coefficient fields. Journal of Number Theory, 18(3):360–370, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900696.

Warkentin:1984:NBL

[302] Peter Warkentin. Über Normformungleichungen mit beschränkter Lösungsmenge. (German) [On norm inequalities with limited solution sets]. Journal of Number Theory, 18(3):371–390, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900702.

Cohen:1984:QRC

[303] Stephen D. Cohen. Quadratic residues and character sums over fields of square order. Journal of Number Theory, 18(3):391–395, June 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900714.

Anonymous:1984:AIVa

[304] Anonymous. Author index for volume 18. Journal of Number Theory, 18(3):396–397, June 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900726.

Anonymous:1984:PAb

[305] Anonymous. Pages 1–130 (August 1984). Journal of Number Theory, 19 (1):??, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1984:EBb

[306] Anonymous. Editorial Board. Journal of Number Theory, 19(1):i, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900878.

Fouche:1984:APH

[307] W. L. Fouché. Arithmetic properties of Heilbronn sums. Journal of Number Theory, 19(1):1–6, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X8490088X.

McCurley:1984:EZF

[308] Kevin S. McCurley. Explicit zero-free regions for Dirichlet L-functions. Journal of Number Theory, 19(1):7–32, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84900891.

Korte:1984:CND

[309] Ulrike Korte. Class numbers of definite binary quadratic lattices over algebraic function fields. Journal of Number Theory, 19(1):33–39, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900908.

Mines:1984:VTC

[310] Ray Mines and Fred Richman. Valuation theory: a constructive view. Journal of Number Theory, 19(1):40–62, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8490091X.

Schmidt:1984:SCA

[311] Wolfgang M. Schmidt. The solubility of certain p-adic equations. Journal of Number Theory, 19(1):63–80, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84900921.

Flahive:1984:ACW

[312] Mary Flahive and Douglas Hensley. Approximation to complex ξ from within a wedge. Journal of Number Theory, 19(1):81–84, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900933.

Apostol:1984:DSR

[313] Tom M. Apostol and Thiennu H. Vu. Dirichlet series related to the Riemann zeta function. Journal of Number Theory, 19(1):85–102, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900945.

Milies:1984:TUI

[314] César Polcino Milies and Sudarshan K. Sehgal. Torsion units in integral group rings of metacyclic groups. Journal of Number Theory, 19(1):103– 114, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X84900957.

Hao:1984:FEQ

[315] Fred H. Hao and Charles J. Parry. The Fermat equation over quadratic fields. Journal of Number Theory, 19(1):115–130, August 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900969.

Anonymous:1984:PO

[316] Anonymous. Pages 131–296 (October 1984). Journal of Number Theory, 19(2):??, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Cusick:1984:VOP

[317] T. W. Cusick and Carl Pomerance. View-obstruction problems, III. Journal of Number Theory, 19(2):131–139, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84900970.

Nakagoshi:1984:NCG

[318] Norikata Nakagoshi. A note on l-class groups of certain algebraic number fields. Journal of Number Theory, 19(2):140–147, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900982.

Goldstein:1984:TZF

[319] Larry Joel Goldstein and Michael J. Razar. The theory of zeta functions of several complex variables, I. Journal of Number Theory, 19(2):148–175, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X84900994.

McCurley:1984:ESC

[320] Kevin S. McCurley. An effective seven cube theorem. Journal of Number Theory, 19(2):176–183, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X84901008.

Rishi:1984:ECC

[321] Dharam Bir Rishi, J. C. Parnami, and A. R. Rajwade. Evaluation of a cubic character sum using the \sqrt{−19} division points of the curve Y^{2} = X^{3} − 2^{3} · 19X + 2 · 19^{2}. Journal of Number Theory, 19(2):184–194, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X8490101X.

Freiling:1984:SNG

[322] Chris Freiling. Some new games and badly approximable numbers. Journal of Number Theory, 19(2):195–202, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84901021.

Tzanakis:1984:CSI

[323] Nicholas Tzanakis. The complete solution in integers of x^{3} + 3y^{3} = 2^{n}. Journal of Number Theory, 19(2):203–208, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84901033.

Spackman:1984:LRS

[324] Kenneth W. Spackman. Linearly recurring solution sequences for equations over finite fields. Journal of Number Theory, 19(2):209–218, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84901045.

Halter-Koch:1984:RKQ

[325] Franz Halter-Koch. Über den 4-rank der Klassengruppe quadratischer Zahlkörper. (German) [On the 4 rank of the class group of quadratic number fields]. Journal of Number Theory, 19(2):219–227, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84901057.

Scarowsky:1984:FAB

[326] M. Scarowsky. On a formal analogue of the Bernoulli numbers. Journal of Number Theory, 19(2):228–232, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84901069.

Vose:1984:ICD

[327] Michael D. Vose. Integers with consecutive divisors in small ratio. Journal of Number Theory, 19(2):233–238, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X84901070.

Gerth:1984:ACS

[328] Frank Gerth and S. W. Graham. Application of a character sum estimate to a 2-class number density. Journal of Number Theory, 19(2):239–247, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X84901082.

Bundschuh:1984:GRI

[329] Peter Bundschuh. Generalization of a recent irrationality result of Mahler. Journal of Number Theory, 19(2):248–253, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84901094.

Nowak:1984:ACC

[330] Werner Georg Nowak. An analogue to a conjecture of S. Chowla and H. Walum. Journal of Number Theory, 19(2):254–262, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84901100.

Muskat:1984:SRP

[331] Joseph B. Muskat. On simultaneous representations of primes by binary quadratic forms. Journal of Number Theory, 19(2):263–282, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84901112.

deOrozco:1984:TGF

[332] Maria Acosta de Orozco and William Yslas Velez. The torsion group of a field defined by radicals. Journal of Number Theory, 19(2):283–294, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X84901124.

Brillhart:1984:LE

[333] John Brillhart. Letter to the editor. Journal of Number Theory, 19(2): 295, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X84901136.

Anonymous:1984:C

[334] Anonymous. Corrigendum. Journal of Number Theory, 19(2):296, October 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84901148.

Anonymous:1984:PD

[335] Anonymous. Pages 297–453 (December 1984). Journal of Number Theory, 19(3):??, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Abramov:1984:BSD

[336] S. A. Abramov and M. J. Rosenbloom. On the Birch–Swinnerton–Dyer conjecture modp. Journal of Number Theory, 19(3):297–300, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900738.

Bateman:1984:PIE

[337] Paul T. Bateman and Emil Grosswald. Positive integers expressible as a sum of three squares in essentially only one way. Journal of Number Theory, 19(3):301–308, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X8490074X.

Buell:1984:ETC

[338] Duncan A. Buell, Richard H. Hudson, and Kenneth S. Williams. Extension of a theorem of Cauchy and Jacobi. Journal of Number Theory, 19(3): 309–340, December 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900751.

Moy:1984:IOS

[339] Allen Moy. The irreducible orthogonal and symplectic Galois representations of a p-adic field (the tame case). Journal of Number Theory, 19 (3):341–344, December 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900763.

Lewin:1984:ISD

[340] L. Lewin. The inner structure of the dilogarithm in algebraic fields. Journal of Number Theory, 19(3):345–373, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900775.

Nagasaka:1984:GDS

[341] Chiaki Nagasaka. On generalized Dedekind sums attached to Dirichlet characters. Journal of Number Theory, 19(3):374–383, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900787.

Emsalem:1984:ILL

[342] M. Emsalem, H. H. Kisilevsky, and D. B. Wales. Indépendance linéaire sur Q de logarithmes P -adiques de nombres algébriques et rang P -adique du groupe des unités d’un corps de nombres. (French) [Linear independence over Q of P -adic logarithms of algebraic numbers and P -adic rank of the group of units in a number field]. Journal of Number Theory, 19 (3):384–391, December 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900799.

Heider:1984:CP

[343] Franz-Peter Heider and Paulgerd Kolvenbach. The construction of SL(2, 3)-polynomials. Journal of Number Theory, 19(3):392–411, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900805.

Ash:1984:CCC

[344] Avner Ash, Daniel Grayson, and Philip Green. Computations of cuspidal cohomology of congruence subgroups of SL(3, Z). Journal of Number Theory, 19(3):412–436, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X84900817.

Silverman:1984:IRR

[345] Joseph H. Silverman. An inequality relating the regulator and the discriminant of a number field. Journal of Number Theory, 19(3):437–442, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X84900829.

Knuth:1984:DCF

[346] Donald E. Knuth. The distribution of continued fraction approximations. Journal of Number Theory, 19(3):443–448, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X84900830.

Anonymous:1984:CGF

[347] Anonymous. “Construction of genus field and some applications” by M. Bhaskaran. Journal of Number Theory, 19(3):449–451, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X84900842. See [1].

Anonymous:1984:TOA

[348] Anonymous. “Transcendence order over Q_{p} in C_{p}” by Alain Escassut. Journal of Number Theory, 19(3):451, December 1984. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/0022314X84900854.

See [223].

Anonymous:1984:AIVb

[349] Anonymous. Author index for volume 19. Journal of Number Theory, 19 (3):452–453, December 1984. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X84900866.

Anonymous:1985:PF

[350] Anonymous. Pages 1–102 (February 1985). Journal of Number Theory, 20(1):??, February 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1985:EBa

[351] Anonymous. Editorial Board. Journal of Number Theory, 20(1):i, February 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900150.

Yu:1985:LFE

[352] Kunrui Yu. Linear forms in elliptic logarithms. Journal of Number Theory, 20(1):1–69, February 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900162.

Sarkozy:1985:DBC

[353] A. Sárközy. On divisors of binomial coefficients, I. Journal of Number Theory, 20(1):70–80, February 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X85900174.

Warshauer:1985:CLN

[354] Max L. Warshauer. Canonical localizers and nonmaximal orders in the Witt setting. Journal of Number Theory, 20(1):81–91, February 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900186.

Zucker:1985:SRS

[355] I. J. Zucker. On the series ∑_{k=1}^{∞}(k^{2k})^{−1}k^{−n} and related sums. Journal of Number Theory, 20(1):92–102, February 1985. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X85900198.

Anonymous:1985:PAa

[356] Anonymous. Pages 103–254 (April 1985). Journal of Number Theory, 20 (2):??, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kanemitsu:1985:CCCa

[357] S. Kanemitsu and R. Sita Rama Chandra Rao. On a conjecture of S. Chowla and of S. Chowla and H. Walum, II. Journal of Number Theory, 20(2):103–120, April 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900319.

Friedberg:1985:TFT

[358] Solomon Friedberg. Theta function transformation formulas and the Weil representation. Journal of Number Theory, 20(2):121–127, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900320.

Thaine:1985:FCF

[359] Francisco Thaine. On the first case of Fermat’s last theorem. Journal of Number Theory, 20(2):128–142, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X85900332.

Prasad:1985:MLC

[360] K. C. Prasad. On a Markoff-like chain. Journal of Number Theory, 20 (2):143–148, April 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900344.

Jaulent:1985:IAN

[361] Jean-François Jaulent. Sur l’indépendance l-adique de nombres algébriques. (French) [On the l-adic independence of algebraic numbers]. Journal of Number Theory, 20(2):149–158, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900356.

Wolfskill:1985:NCR

[362] J. Wolfskill. A note concerning reduced algebraic numbers and equivalence of binary forms. Journal of Number Theory, 20(2):159–161, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900368.

Wolfskill:1985:DRN

[363] J. Wolfskill. The distribution of reduced numbers in an ideal of a real cubic number field. Journal of Number Theory, 20(2):162–171, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8590037X.

Krass:1985:EDD

[364] S. Krass. Estimates for n-dimensional Diophantine approximation constants for n > −4. Journal of Number Theory, 20(2):172–176, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900381.

Buchmann:1985:GVUa

[365] Johannes Buchmann. A generalization of Voronoi’s unit algorithm I. Journal of Number Theory, 20(2):177–191, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X85900393.

Buchmann:1985:GVUb

[366] Johannes Buchmann. A generalization of Voronoi’s unit algorithm II. Journal of Number Theory, 20(2):192–209, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8590040X.

Sagan:1985:CAG

[367] Bruce E. Sagan. Congruences via abelian groups. Journal of Number Theory, 20(2):210–237, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X85900411.

Horie:1985:ISI

[368] Kuniaki Horie. On the index of the Stickelberger ideal and the cyclotomic regulator. Journal of Number Theory, 20(2):238–253, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900423.

Anonymous:1985:Aa

[369] Anonymous. Announcement. Journal of Number Theory, 20(2):254, April 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900435.

Anonymous:1985:PJ

[370] Anonymous. Pages 255–406 (June 1985). Journal of Number Theory, 20 (3):??, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kanemitsu:1985:CCCb

[371] S. Kanemitsu and R. Sita Rama Chandra Rao. On a conjecture of S. Chowla and of S. Chowla and H. Walum, I. Journal of Number Theory, 20(3):255–261, June 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900204.

Massell:1985:RCF

[372] Paul B. Massell. Ray class field extensions of real quadratic fields and solvability of congruences. Journal of Number Theory, 20(3):262–272, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900216.

Soicher:1985:CGG

[373] Leonard Soicher and John McKay. Computing Galois groups over the rationals. Journal of Number Theory, 20(3):273–281, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900228.

Benz:1985:VPG

[374] H. Benz and H. Zassenhaus. Über verschränkte Produktordnungen. (German) [On restricted product rules]. Journal of Number Theory, 20(3):282– 298, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X8590023X.

Massy:1985:SEP

[375] Richard Massy. Solutions explicites de problèmes de plongement. (French) [Explicit solutions of embedding problems]. Journal of Number Theory, 20(3):299–314, June 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900241.

Scourfield:1985:UCR

[376] E. J. Scourfield. A uniform coprimality result for some arithmetic functions. Journal of Number Theory, 20(3):315–353, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900253.

Warlimont:1985:SNN

[377] Richard Warlimont. On the set of natural numbers which only yield orders of Abelian groups. Journal of Number Theory, 20(3):354–362, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900265.

Nash:1985:CSA

[378] John C. M. Nash and Melvyn B. Nathanson. Cofinite subsets of asymptotic bases for the positive integers. Journal of Number Theory, 20(3):363– 372, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X85900277.

Hlawka:1985:CLS

[379] Edmund Hlawka. Carl Ludwig Siegel (1896–1981). Journal of Number Theory, 20(3):373–404, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X85900289.

Anonymous:1985:Ab

[380] Anonymous. Announcement. Journal of Number Theory, 20(3):405, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900290.

Anonymous:1985:AIVa

[381] Anonymous. Author index for volume 20. Journal of Number Theory, 20 (3):406, June 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X85900307.

Anonymous:1985:PAb

[382] Anonymous. Pages 1–120 (August 1985). Journal of Number Theory, 21 (1):??, August 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1985:EBb

[383] Anonymous. Editorial Board. Journal of Number Theory, 21(1):i, August 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900083.

Turner:1985:NTF

[384] Stuart Turner. Nonarchimedean theta functions and the Néron symbol. Journal of Number Theory, 21(1):1–16, August 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900095.

Boyd:1985:PNN

[385] David W. Boyd. Pisot numbers in the neighbourhood of a limit point, I. Journal of Number Theory, 21(1):17–43, August 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900101.

Saito:1985:CFT

[386] Shuji Saito. Class field theory for curves over local fields. Journal of Number Theory, 21(1):44–80, August 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X85900113.

Spiro:1985:HON

[387] Claudia Spiro. How often is the number of divisors of n a divisor of n? Journal of Number Theory, 21(1):81–100, August 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900125.

Joris:1985:GCD

[388] H. Joris, C. Oestreicher, and J. Steinig. The greatest common divisor of certain sets of binomial coefficients. Journal of Number Theory, 21 (1):101–119, August 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900137.

Anonymous:1985:Ac

[389] Anonymous. Announcement. Journal of Number Theory, 21(1):120, August 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900149.

Anonymous:1985:PO

[390] Anonymous. Pages 121–244 (October 1985). Journal of Number Theory, 21(2):??, October 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Goncalves:1985:FSG

[391] Jairo Zacarias Gonçalves. Free subgroups in the group of units of group rings II. Journal of Number Theory, 21(2):121–127, October 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900447.

Shallit:1985:IPA

[392] J. O. Shallit. On infinite products associated with sums of digits. Journal of Number Theory, 21(2):128–134, October 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X85900459.

Gekeler:1985:PED

[393] Ernst-Ulrich Gekeler. A product expansion for the discriminant function of Drinfeld modules of rank two. Journal of Number Theory, 21(2):135– 140, October 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X85900460.

Beukers:1985:SCA

[394] F. Beukers. Some congruences for the Apery numbers. Journal of Number Theory, 21(2):141–155, October 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X85900472.

Kirschenhofer:1985:DFG

[395] Peter Kirschenhofer and Robert F. Tichy. Zur diskrepanz von 0, 1-Folgen. (German) [On the discrepancy of 0, 1 sequences]. Journal of Number Theory, 21(2):156–175, October 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X85900484.

Traina:1985:CPM

[396] Charles Traina. The conjugacy problem of the modular group and the class number of real quadratic number fields. Journal of Number Theory, 21(2):176–184, October 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900496.

Li:1985:ICN

[397] Wen-Ch’ing Winnie Li and B. Z. Moroz. On ideal classes of number fields containing integral ideals of equal norms. Journal of Number Theory, 21 (2):185–203, October 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900502.

Kramer:1985:RPC

[398] Kenneth Kramer. Residue properties of certain quadratic units. Journal of Number Theory, 21(2):204–213, October 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X85900514.

Abouzahra:1985:PAN

[399] M. Abouzahra and L. Lewin. The polylogarithm in algebraic number fields. Journal of Number Theory, 21(2):214–244, October 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900526.

Anonymous:1985:PD

[400] Anonymous. Pages 245–413 (December 1985). Journal of Number Theory, 21(3):??, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Meyer:1985:CCW

[401] Jacques Meyer. Une conjecture de Chowla et Walum. (French) [A conjecture of Chowla and Walum]. Journal of Number Theory, 21(3):245–255, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X85900538.

Bhaskaran:1985:SFR

[402] M. Bhaskaran. Some further remarks on genus field. Journal of Number Theory, 21(3):256–259, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X8590054X.

Cornell:1985:CNC

[403] Gary Cornell and Lawrence C. Washington. Class numbers of cyclotomic fields. Journal of Number Theory, 21(3):260–274, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X85900551.

Vulakh:1985:MRI

[404] L. Ya Vulakh. On minima of rational indefinite quadratic forms. Journal of Number Theory, 21(3):275–285, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X85900563.

Hensley:1985:NPI

[405] Douglas Hensley. The number of positive integers ≤ x and free of prime factors > y. Journal of Number Theory, 21(3):286–298, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/

0022314X85900575. See comment [435].

Lloyd-Smith:1985:MDA

[406] C. W. Lloyd-Smith. On minimal diameters of algebraic integers in J-fields. Journal of Number Theory, 21(3):299–318, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900587.

Erdos:1985:LRV

[407] P. Erdős, A. Sárközy, and C. Pomerance. On locally repeated values of certain arithmetic functions, I. Journal of Number Theory, 21(3):319–332, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X85900599.

Sarnak:1985:CNI

[408] Peter C. Sarnak. Class numbers of indefinite binary quadratic forms II. Journal of Number Theory, 21(3):333–346, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X85900605.

Osgood:1985:SET

[409] Charles F. Osgood. Sometimes effective Thue–Siegel–Roth–Schmidt– Nevanlinna bounds, or better. Journal of Number Theory, 21(3):347–389, December 1985. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X85900617.

Beauzamy:1985:EPP

[410] B. Beauzamy and P. Enflo. Estimations de produits de polynômes. (French) [Polynomial product estimates]. Journal of Number Theory, 21 (3):390–412, December 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900629.

Anonymous:1985:AIVb

[411] Anonymous. Author index for volume 21. Journal of Number Theory, 21(3):413, December 1985. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X85900630.

Anonymous:1986:PJa

[412] Anonymous. Pages 1–123 (January 1986). Journal of Number Theory, 22 (1):??, January 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1986:EBa

[413] Anonymous. Editorial Board. Journal of Number Theory, 22(1):i, January 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900259.

Chowla:1986:LPQ

[414] S. Chowla, J. Cowles, and M. Cowles. The least prime quadratic residue and the class number. Journal of Number Theory, 22(1):1–3, January 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900260.

Faure:1986:LBP

[415] Henri Faure. Lemme de Bohl pour les suites de Van Der Corput generalisées. (French) [Bohl’s lemma for the generalized sequences of Van Der Corput]. Journal of Number Theory, 22(1):4–20, January 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900272.

Monzingo:1986:EEJ

[416] M. G. Monzingo. An elementary evaluation of the Jacobsthal sum. Journal of Number Theory, 22(1):21–25, January 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900284.

Volklein:1986:FLG

[417] Helmut Völklein. Fahnentransitive lineare Gruppen über globalen Körpern und über absolut-algebraischen Körpern positiver charakteristik. (German) [Flag-transitive linear groups over global bodies and over absolutely algebraic bodies with positive characteristics]. Journal of Number Theory, 22(1):26–45, January 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X86900296.

Hensley:1986:PCF

[418] Douglas Hensley. A property of the counting function of integers with no large prime factors. Journal of Number Theory, 22(1):46–74, January 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900302.

Deshommes:1986:CRA

[419] Bernadette Deshommes. Critères de rationalité et application à la série génératrice d’un système d’équations à coefficients dans un corps local. (French) [Rationality criteria and application to the generating series of a system of coefficient equations in a local field]. Journal of Number Theory, 22(1):75–114, January 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X86900314.

Nakajima:1986:GMS

[420] Shōichi Nakajima. Galois module structure of cohomology groups for tamely ramified coverings of algebraic varieties. Journal of Number Theory, 22(1):115–123, January 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X86900326.

Anonymous:1986:PF

[421] Anonymous. Pages 125–248 (February 1986). Journal of Number Theory, 22(2):??, February 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Skula:1986:NII

[422] Ladislav Skula. A note on the index of irregularity. Journal of Number Theory, 22(2):125–138, February 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900648.

Coquet:1986:SFI

[423] Jean Coquet and Pascal Van Den Bosch. A summation formula involving Fibonacci digits. Journal of Number Theory, 22(2):139–146, February 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8690065X.

Mahler:1986:SMG

[424] Kurt Mahler. The successive minima in the geometry of numbers and the distinction between algebraic and transcendental numbers. Journal of Number Theory, 22(2):147–160, February 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900661.

Coquet:1986:PSD

[425] Jean Coquet. Power sums of digital sums. Journal of Number Theory, 22 (2):161–176, February 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X86900673.

Ruck:1986:CGS

[426] H.-G. Rück. Class groups and L-series of function fields. Journal of Number Theory, 22(2):177–189, February 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900685.

Mason:1986:NFE

[427] R. C. Mason. Norm form equations I. Journal of Number Theory, 22 (2):190–207, February 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X86900697.

Grupp:1986:FLS

[428] F. Grupp and H.-E. Richert. The functions of the linear sieve. Journal of Number Theory, 22(2):208–239, February 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900703.

Feng:1986:NIP

[429] Keqin Feng. A note on irregular prime polynomials in cyclotomic function field theory. Journal of Number Theory, 22(2):240–245, February 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900715.

Godwin:1986:PSC

[430] H. J. Godwin. A parametrized set of cyclic cubic fields with even class-number. Journal of Number Theory, 22(2):246–248, February 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900727.

Anonymous:1986:PMa

[431] Anonymous. Pages 249–350 (March 1986). Journal of Number Theory, 22(3):??, March 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Halter-Koch:1986:BQF

[432] Franz Halter-Koch. Binäre quadratische Formen und rationale Zerlegungsgesetze. I. (German) [Binary quadratic forms and rational decomposition laws. I]. Journal of Number Theory, 22(3):249–270, March 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900107.

Kalpazidou:1986:SAR

[433] Sofia Kalpazidou. Some asymptotic results on digits of the nearest integer continued fraction. Journal of Number Theory, 22(3):271–279, March 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900119.

Bantle:1986:PES

[434] G. Bantle and F. Grupp. On a problem of Erdős and Szemerédi. Journal of Number Theory, 22(3):280–288, March 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900120.

Hildebrand:1986:NPI

[435] Adolf Hildebrand. On the number of positive integers ≤ x and free of prime factors > y. Journal of Number Theory, 22(3):289–307, March 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900132. See [405].

Hall:1986:EMD

[436] R. R. Hall and G. Tenenbaum. Les ensembles de multiples et la densité divisorielle. (French) [Sets of multiples and divisor density]. Journal of Number Theory, 22(3):308–333, March 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900144.

Hinz:1986:NAC

[437] Jürgen G. Hinz. A note on Artin’s conjecture in algebraic number fields. Journal of Number Theory, 22(3):334–349, March 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900156.

Anonymous:1986:AIVa

[438] Anonymous. Author index for volume 22. Journal of Number Theory, 22(3):350, March 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X86900168.

Anonymous:1986:PMb

[439] Anonymous. Pages 1–148 (May 1986). Journal of Number Theory, 23 (1):??, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1986:EBb

[440] Anonymous. Editorial Board. Journal of Number Theory, 23(1):i, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900016.

Cassou-Nogues:1986:SDI

[441] Pierrette Cassou-Nogués. Séries de Dirichlet et intégrales associées à un polynôme à deux indéterminées. (French) [Dirichlet series and integrals associated with a polynomial with two indeterminates]. Journal of Number Theory, 23(1):1–54, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X86900028.

Wu:1986:PCF

[442] Trong Wu. On the proof of continued fraction expansions for irrationals. Journal of Number Theory, 23(1):55–59, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8690003X.

Tubbs:1986:MAI

[443] Robert Tubbs. On the measure of algebraic independence of certain values of elliptic functions. Journal of Number Theory, 23(1):60–79, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900041.

Wajngurt:1986:RSD

[444] Clara Wajngurt. Rational solutions of diophantine equations isomorphic to elliptic curves with applications to complex multiplication. Journal of Number Theory, 23(1):80–85, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900053.

Hikita:1986:CCN

[445] Mitsuhiko Hikita. On the congruences for the class numbers of the quadratic fields whose discriminants are divisible by 8. Journal of Number Theory, 23(1):86–101, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X86900065.

Hashimoto:1986:IMT

[446] Ki ichiro Hashimoto and Robert J. Sibner. Involutive modular transformations on the Siegel upper half plane and an application to representations of quadratic forms. Journal of Number Theory, 23(1):102–110, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900077.

Driencourt:1986:PAC

[447] Yves Driencourt. Une présentation adélique de certaines fonctions analogues A log |η(z)|. (French) [An adic presentation of some analogous functions A log |η(z)|]. Journal of Number Theory, 23(1):111–135, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900089.

Greene:1986:CSE

[448] J. Greene and D. Stanton. A character sum evaluation and Gaussian hypergeometric series. Journal of Number Theory, 23(1):136–148, May 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900090.

Anonymous:1986:PJb

[449] Anonymous. Pages 149–278 (June 1986). Journal of Number Theory, 23 (2):??, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Sitaramachandrarao:1986:S

[450] R. Sitaramachandrarao and P. V. Krishnaiah. On the sums ∑_{n≤x} A(f(n)) and ∑_{p≤x} A(f(p)). Journal of Number Theory, 23(2):149–168, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900867.

Kaplan:1986:PEC

[451] Pierre Kaplan and Kenneth S. Williams. Pell’s equations X^{2} − mY^{2} = −1, −4 and continued fractions. Journal of Number Theory, 23(2):169– 182, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X86900879.

Pintz:1986:SCG

[452] Jànos Pintz and Saverio Salerno. Some consequences of the general Riemann hypothesis in the comparative theory of primes. Journal of Number Theory, 23(2):183–194, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X86900880.

Matzat:1986:RMG

[453] B. Heinrich Matzat and Andreas Zeh-Marschke. Realisierung der Mathieugruppen M_{11} und M_{12} als Galoisgruppen über Q. (German) [Realization of the Mathieu groups M_{11} and M_{12} as Galois groups over Q]. Journal of Number Theory, 23(2):195–202, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900892.

Shemanske:1986:TQF

[454] Thomas R. Shemanske. Ternary quadratic forms and quaternion algebras. Journal of Number Theory, 23(2):203–209, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900909.

Frankl:1986:UFF

[455] P. Frankl and Z. Füredi. Union-free families of sets and equations over fields. Journal of Number Theory, 23(2):210–218, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900910.

Tzanakis:1986:DE

[456] Nikos Tzanakis and John Wolfskill. On the diophantine equation y^{2} = 4q^{n} + 4q + 1. Journal of Number Theory, 23(2):219–237, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900922.

Fukuda:1986:IER

[457] Takashi Fukuda and Keiichi Komatsu. On the λ invariants of Z_{p}- extensions of real quadratic fields. Journal of Number Theory, 23(2):238– 242, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X86900934.

Smyth:1986:AMR

[458] C. J. Smyth. Additive and multiplicative relations connecting conjugate algebraic numbers. Journal of Number Theory, 23(2):243–254, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900946.

Friedberg:1986:TFA

[459] Solomon Friedberg. On theta functions associated to indefinite quadratic forms. Journal of Number Theory, 23(2):255–267, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900958.

Pavone:1986:RTS

[460] Marco Pavone. A remark on a theorem of Serret. Journal of Number Theory, 23(2):268–278, June 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X8690096X.

Anonymous:1986:PJc

[461] Anonymous. Pages 279–406 (July 1986). Journal of Number Theory, 23 (3):??, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Lampert:1986:AAE

[462] David Lampert. Algebraic p-adic expansions. Journal of Number Theory, 23(3):279–284, July 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X86900739.

Andrews:1986:N

[463] George E. Andrews. EΥPHKA! num = Δ + Δ + Δ. Journal of Number Theory, 23(3):285–293, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X86900740.

Satge:1986:GSC

[464] Philippe Satge. Groupes de Selmer et corps cubiques. (French) [Selmer groups and cubic bodies]. Journal of Number Theory, 23(3):294–317, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900752.

Xianke:1986:CCA

[465] Zhang Xianke. Counterexample and correction about genus fields of number fields. Journal of Number Theory, 23(3):318–321, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900764.

Gras:1986:RNF

[466] Georges Gras. Remarks on K_{2} of number fields. Journal of Number Theory, 23(3):322–335, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X86900776.

Cougnard:1986:BNR

[467] Jean Cougnard. Bases normales relatives dans certaines extensions cyclotomiques. (French) [Relative normal bases in certain cyclotomic extensions]. Journal of Number Theory, 23(3):336–346, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900788.

Gras:1986:NMA

[468] Marie-Nicole Gras. Non monogénéité de l’anneau des entiers des extensions cycliques de Q de degré premier l ≥ 5. (French) [Nonuniformity of the ring of integers of cyclic extensions of Q of prime degree l ≥ 5]. Journal of Number Theory, 23(3):347–353, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X8690079X.

Nishioka:1986:AIC

[469] Kumiko Nishioka. Algebraic independence of certain power series of algebraic numbers. Journal of Number Theory, 23(3):354–364, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900806.

MacLeod:1986:ERD

[470] R. A. MacLeod. An extremal result for divisor functions. Journal of Number Theory, 23(3):365–366, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900818.

Deanin:1986:PAC

[471] Alice A. Deanin. Periodicity of P -adic continued fraction expansions. Journal of Number Theory, 23(3):367–387, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8690082X.

Mills:1986:CFC

[472] W. H. Mills and David P. Robbins. Continued fractions for certain algebraic power series. Journal of Number Theory, 23(3):388–404, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900831.

Girstmair:1986:LE

[473] Kurt Girstmair. Letter to the editor. Journal of Number Theory, 23(3): 405, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X86900843.

Anonymous:1986:AIVb

[474] Anonymous. Author index for volume 23. Journal of Number Theory, 23 (3):406, July 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X86900855.

Anonymous:1986:PS

[475] Anonymous. Pages 1–125 (September 1986). Journal of Number Theory, 24(1):??, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1986:EBc

[476] Anonymous. Editorial Board. Journal of Number Theory, 24(1):i, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8690051X.

Hilliker:1986:ULC

[477] David Lee Hilliker and E. G. Straus. Uniqueness of linear combinations ( mod p). Journal of Number Theory, 24(1):1–6, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900521.

Mollin:1986:DEC

[478] R. A. Mollin. Diophantine equations and class numbers. Journal of Number Theory, 24(1):7–19, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900533.

Mahler:1986:NTP

[479] Kurt Mahler. A new transfer principle in the geometry of numbers. Journal of Number Theory, 24(1):20–34, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900545.

Hlawka:1986:GSM

[480] Edmund Hlawka. Gleichverteilung und ein Satz von Müntz. (German) [Equal distribution and a sentence by Müntz]. Journal of Number Theory, 24(1):35–46, September 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X86900557.

Nemes:1986:PVL

[481] István Nemes and Attila Pethö. Polynomial values in linear recurrences, II. Journal of Number Theory, 24(1):47–53, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900569.

Schoissengeier:1986:DI

[482] Johannes Schoissengeier. On the discrepancy of (nα), II. Journal of Number Theory, 24(1):54–64, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900570.

Tijdeman:1986:ARM

[483] R. Tijdeman. Approximation of real matrices by integral matrices. Journal of Number Theory, 24(1):65–69, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900582.

deWeger:1986:ALA

[484] B. M. M. de Weger. Approximation lattices of p-adic numbers. Journal of Number Theory, 24(1):70–88, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900594.

Yokota:1986:CMB

[485] Hisashi Yokota. On a conjecture of M. N. Bleicher and P. Erdős. Journal of Number Theory, 24(1):89–94, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900600.

Apostol:1986:NHP

[486] Tom M. Apostol. On the nonvanishing of homogeneous product sums. Journal of Number Theory, 24(1):95–106, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900612.

Federer:1986:NII

[487] Leslie Jane Federer. A note on Iwasawa invariants and the main conjecture. Journal of Number Theory, 24(1):107–113, September 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900624.

Kalpazidou:1986:NCF

[488] Sofia Kalpazidou. On nearest continued fractions with stochastically independent and identically distributed digits. Journal of Number Theory, 24 (1):114–125, September 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X86900636.

Anonymous:1986:PO

[489] Anonymous. Pages 127–248 (October 1986). Journal of Number Theory, 24(2):??, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Wallace:1986:EFH

[490] D. I. Wallace. Explicit form of the hyperbolic term in the Selberg trace formula for Sl(3, Z) and Pell’s equation for hyperbolics in Sl(3, Z). Journal of Number Theory, 24(2):127–133, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900971.

Bantle:1986:DLS

[491] G. Bantle. On the A-dependence of the linear sieve with application to polynomial sequences. Journal of Number Theory, 24(2):134–153, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900983.

Grupp:1986:DDE

[492] F. Grupp. On difference-differential equations in the theory of sieves. Journal of Number Theory, 24(2):154–173, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900995.

Nagasaka:1986:EVD

[493] Chiaki Nagasaka. Exceptional values of the Dedekind symbol. Journal of Number Theory, 24(2):174–180, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86901009.

Mozzochi:1986:DBC

[494] C. J. Mozzochi. On the difference between consecutive primes. Journal of Number Theory, 24(2):181–187, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86901010.

Chowla:1986:D

[495] S. Chowla, B. Dwork, and Ronald Evans. On the modp^{2} determination of \binom{(p−1)/2}{(p−1)/4}. Journal of Number Theory, 24(2):188–196, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86901022.

Niederreiter:1986:ITM

[496] Harald Niederreiter. On an irrationality theorem of Mahler and Bundschuh. Journal of Number Theory, 24(2):197–199, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86901034.

Lagarias:1986:SDA

[497] Jeffrey C. Lagarias and Johan T. Hastad. Simultaneous Diophantine approximation of rationals by rationals. Journal of Number Theory, 24(2): 200–228, October 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X86901046.

Kato:1986:DFA

[498] Kazuya Kato and Takako Kuzumaki. The dimension of fields and algebraic K-theory. Journal of Number Theory, 24(2):229–244, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86901058.

Ribenboim:1986:NCE

[499] Paulo Ribenboim. A note on Catalan’s equation. Journal of Number Theory, 24(2):245–248, October 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X8690106X.

Anonymous:1986:PN

[500] Anonymous. Pages 249–374 (November 1986). Journal of Number Theory, 24(3):??, November 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Yokota:1986:LDE

[501] Hisashi Yokota. Length and denominators of Egyptian fractions. Journal of Number Theory, 24(3):249–258, November 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X86900338.

Brown:1986:DTN

[502] Gavin Brown, William Moran, and Charles E. M. Pearce. A decomposition theorem for numbers in which the summands have prescribed normality properties. Journal of Number Theory, 24(3):259–271, November 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8690034X.

Moroz:1986:IPN

[503] B. Z. Moroz. Integral points on norm-form varieties. Journal of Number Theory, 24(3):272–283, November 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900351.

Rhin:1986:APS

[504] Georges Rhin and Philippe Toffin. Approximants de Padé simultanés de logarithmes. (French) [Simultaneous Padé approximants of logarithms]. Journal of Number Theory, 24(3):284–297, November 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X86900363.

Kovacs:1986:CAF

[505] Katalin Kovács. On the characterization of additive functions with monotonic norm. Journal of Number Theory, 24(3):298–304, November 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X86900375.

Bruen:1986:PFG

[506] Aiden A. Bruen, Christian U. Jensen, and Noriko Yui. Polynomials with Frobenius groups of prime degree as Galois groups II. Journal of Number Theory, 24(3):305–359, November 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X86900387.

Gow:1986:CSW

[507] R. Gow. Construction of some wreath products as Galois groups of normal real extensions of the rationals. Journal of Number Theory, 24(3):360–372, November 1986. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X86900399.

Anonymous:1986:AIVc

[508] Anonymous. Author index for volume 24. Journal of Number Theory, 24 (3):373–374, November 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X86900405.

Anonymous:1987:PJa

[509] Anonymous. Pages 1–131 (January 1987). Journal of Number Theory, 25 (1):??, January 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1987:EBa

[510] Anonymous. Editorial Board. Journal of Number Theory, 25(1):i, January 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900114.

Sitaramachandrarao:1987:FR

[511] R. Sitaramachandrarao. A formula of S. Ramanujan. Journal of Number Theory, 25(1):1–19, January 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X87900126.

Schneps:1987:IAF

[512] Leila Schneps. On the μ-invariant of p-adic L-functions attached to elliptic curves with complex multiplication. Journal of Number Theory, 25(1):20– 33, January 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X87900138.

DiazyDiaz:1987:PDC

[513] F. Diaz y Diaz. Petits discriminants des corps de nombres totalement imaginaires de degré 8. (French) [Small discriminants of fields of totally imaginary numbers of degree 8]. Journal of Number Theory, 25(1):34–52, January 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X8790014X.

Saito:1987:EUK

[514] Hiroshi Saito. Elliptic units and a Kummer’s criterion for imaginary quadratic fields. Journal of Number Theory, 25(1):53–71, January 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900151.

Dilcher:1987:ICG

[515] Karl Dilcher. Irreducibility of certain generalized Bernoulli polynomials belonging to quadratic residue characters. Journal of Number Theory, 25(1):72–80, January 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900163.

Hildebrand:1987:NPF

[516] Adolf Hildebrand. On the number of prime factors of integers without large prime divisors. Journal of Number Theory, 25(1):81–106, January 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900175.

Panaitopol:1987:RCI

[517] L. Panaitopol and D. Ştefānescu. A resultant condition for the irreducibility of the polynomials. Journal of Number Theory, 25(1):107–111, January 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900187.

Poulakis:1987:CMS

[518] Dimitrios Poulakis. La courbe modulaire X_{0}(125) et sa jacobienne. (French) [The modular curve X_{0}(125) and its Jacobian]. Journal of Number Theory, 25(1):112–131, January 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900199.

Anonymous:1987:PF

[519] Anonymous. Pages 133–248 (February 1987). Journal of Number Theory, 25(2):??, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Jacquinot:1987:DPE

[520] Olivier Jacquinot and Kenneth A. Ribet. Deficient points on extensions of abelian varieties by G_{m}. Journal of Number Theory, 25(2):133–151, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X87900205.

Ribet:1987:CRF

[521] Kenneth A. Ribet. Cohomological realization of a family of 1-motives. Journal of Number Theory, 25(2):152–161, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900217.

Richert:1987:RCE

[522] Norman Richert. Regular chains for equivalent complex irrationals. Journal of Number Theory, 25(2):162–168, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900229.

Wimmer:1987:RMR

[523] Harald K. Wimmer. Realizations of matrices of rational numbers. Journal of Number Theory, 25(2):169–183, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900230.

Faisant:1987:QRT

[524] Alain Faisant and Georges Philibert. Quelques résultats de transcendance liés à l’invariant modulaire j. (French) [Some results of transcendence linked to the modular invariant j]. Journal of Number Theory, 25(2): 184–200, February 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900242.

Beukers:1987:ACA

[525] F. Beukers. Another congruence for the Apéry numbers. Journal of Number Theory, 25(2):201–210, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900254.

Zun:1987:NIS

[526] Shan Zun. A note on irrationality of some numbers. Journal of Number Theory, 25(2):211–212, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900266.

Madan:1987:RAI

[527] Manohar L. Madan and Horst G. Zimmer. Relations among Iwasawa invariants. Journal of Number Theory, 25(2):213–219, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900278.

Wakabayashi:1987:AVM

[528] Isao Wakabayashi. Algebraic values of meromorphic functions on Riemann surfaces. Journal of Number Theory, 25(2):220–229, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8790028X.

Osada:1987:GGP

[529] Hiroyuki Osada. The Galois groups of the polynomials X^{n} + aX^{1} + b. Journal of Number Theory, 25(2):230–238, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900291.

Mason:1987:NFE

[530] R. C. Mason. Norm form equations. V. Degenerate modules. Journal of Number Theory, 25(2):239–248, February 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900308.

Anonymous:1987:PMa

[531] Anonymous. Pages 249–395 (March 1987). Journal of Number Theory, 25(3):??, March 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Friedlander:1987:DSI

[532] J. B. Friedlander and J. C. Lagarias. On the distribution in short intervals of integers having no large prime factor. Journal of Number Theory, 25 (3):249–273, March 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X8790031X.

Matthews:1987:GSA

[533] K. R. Matthews and G. M. Leigh. A generalization of the Syracuse algorithm in F_{q}[x]. Journal of Number Theory, 25(3):274–278, March 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900321.

Tubbs:1987:AGS

[534] Robert Tubbs. Algebraic groups and small transcendence degree I. Journal of Number Theory, 25(3):279–307, March 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900333.

Nijenhuis:1987:PPF

[535] Albert Nijenhuis and Herbert S. Wilf. Periodicities of partition functions and Stirling numbers modulo p. Journal of Number Theory, 25(3):308– 312, March 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X87900345.

Saito:1987:MTV

[536] Hiroshi Saito. On missing trace values for the eta multipliers. Journal of Number Theory, 25(3):313–327, March 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900357.

Pettet:1987:TTR

[537] Martin R. Pettet and R. Sitaramachandrarao. Three-term relations for Hardy sums. Journal of Number Theory, 25(3):328–339, March 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900369.

Marciniak:1987:TUI

[538] Z. Marciniak, J. Ritter, S. K. Sehgal, and A. Weiss. Torsion units in integral group rings of some metabelian groups, II. Journal of Number Theory, 25(3):340–352, March 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X87900370.

Gurak:1987:RNG

[539] S. Gurak. Relative norms of Gauss sums. Journal of Number Theory, 25(3):353–359, March 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900382.

Erdos:1987:EFG

[540] P. Erdős, M. Ram Murty, and V. Kumar Murty. On the enumeration of finite groups. Journal of Number Theory, 25(3):360–378, March 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900394.

Gillard:1987:TML

[541] R. Gillard. Transformation de Mellin–Leopoldt des fonctions elliptiques. (French) [Mellin–Leopoldt transformation of elliptic functions]. Journal of Number Theory, 25(3):379–393, March 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900400.

Anonymous:1987:AIVa

[542] Anonymous. Author index for volume 25. Journal of Number Theory, 25(3):394–395, March 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900412.

Anonymous:1987:PMb

[543] Anonymous. Pages 1–116 (May 1987). Journal of Number Theory, 26 (1):??, May 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1987:EBb

[544] Anonymous. Editorial Board. Journal of Number Theory, 26(1):i, May 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900904.

Hung:1987:TST

[545] D. C. Hung. Theta series of ternary quadratic forms. Journal of Number Theory, 26(1):1–7, May 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900916.

Buchmann:1987:CUC

[546] Johannes Buchmann. On the computation of units and class numbers by a generalization of Lagrange’s algorithm. Journal of Number Theory, 26(1): 8–30, May 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X87900928.

Buchmann:1987:PLG

[547] Johannes Buchmann. On the period length of the generalized Lagrange algorithm. Journal of Number Theory, 26(1):31–37, May 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8790093X.

Hsia:1987:PSE

[548] J. S. Hsia. On primitive spinor exceptional representations. Journal of Number Theory, 26(1):38–45, May 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900941.

Denef:1987:APS

[549] J. Denef and L. Lipshitz. Algebraic power series and diagonals. Journal of Number Theory, 26(1):46–67, May 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900953.

Tichy:1987:DSS

[550] Robert F. Tichy and Gerhard Turnwald. On the discrepancy of some special sequences. Journal of Number Theory, 26(1):68–78, May 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900965.

Lundell:1987:DGB

[551] Albert T. Lundell. On the denominator of generalized Bernoulli numbers. Journal of Number Theory, 26(1):79–88, May 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900977.

Iwaniec:1987:MHP

[552] Henryk Iwaniec and András Sárközy. On a multiplicative hybrid problem. Journal of Number Theory, 26(1):89–95, May 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900989.

Tzanakis:1987:DEA

[553] Nikos Tzanakis and John Wolfskill. The diophantine equation x^{2} = 4q^{a2}+ 4q + 1, with an application to coding theory. Journal of Number Theory, 26(1):96–116, May 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900990.

Anonymous:1987:PJb

[554] Anonymous. Pages 117–236 (June 1987). Journal of Number Theory, 26 (2):??, June 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Miki:1987:LCA

[555] Hiroo Miki. On the Leopoldt conjecture on the p-adic regulators. Journal of Number Theory, 26(2):117–128, June 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900734.

Grekos:1987:DG

[556] Georges Grekos and Bodo Volkmann. On densities and gaps. Journal of Number Theory, 26(2):129–148, June 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900746.

Guralnick:1987:GMI

[557] Robert M. Guralnick. The genus of a module II. Roiter’s theorem, power cancellation and extension of scalars. Journal of Number Theory, 26 (2):149–165, June 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900758.

Huttner:1987:ICI

[558] M. Huttner. Irrationalité de certaines integrales hypergéométriques. (French) [Irrationality of certain hypergomometric integrals]. Journal of Number Theory, 26(2):166–178, June 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X8790076X.

Hensley:1987:DNF

[559] Doug Hensley. The distribution of the number of factors in a factorization. Journal of Number Theory, 26(2):179–191, June 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900771.

Burde:1987:PAZ

[560] Klaus Burde. Das Problem der Abzählreime und Zahlentwicklungen mit gebrochenen basen. (German) [The problem of counting rhymes and number developments with broken bases]. Journal of Number Theory, 26 (2):192–209, June 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900783.

Feit:1987:ESA

[561] Walter Feit and Paul Vojta. Examples of some Q-admissible groups. Journal of Number Theory, 26(2):210–226, June 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900795.

Kleinert:1987:U

[562] Ernst Kleinert. Units in Z[Q_{p}]. Journal of Number Theory, 26(2):227– 236, June 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X87900801.

Anonymous:1987:PJc

[563] Anonymous. Pages 237–368 (July 1987). Journal of Number Theory, 26 (3):??, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Chidambaraswamy:1987:PVP

[564] J. Chidambaraswamy and R. Sitaramachandrarao. On the probability that the values of m polynomials have a given g.c.d. Journal of Number Theory, 26(3):237–245, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X87900813.

Berend:1987:DDS

[565] Daniel Berend. Dense (mod1) dilated semigroups of algebraic numbers. Journal of Number Theory, 26(3):246–256, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900825.

Delzell:1987:CPN

[566] Charles N. Delzell. Continuous Pythagoras numbers for rational quadratic forms. Journal of Number Theory, 26(3):257–273, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900837.

Yasumoto:1987:HIS

[567] M. Yasumoto. Hilbert irreducibility sequences and nonstandard arithmetic. Journal of Number Theory, 26(3):274–285, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900849.

Vaserstein:1987:WPA

[568] L. N. Vaserstein. Waring’s problem for algebras over fields. Journal of Number Theory, 26(3):286–298, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900850.

Vaserstein:1987:WPC

[569] L. N. Vaserstein. Waring’s problem for commutative rings. Journal of Number Theory, 26(3):299–307, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900862.

Hrycaj:1987:PMF

[570] Jordan Hrycaj. Polynomial mappings of finitely generated Z-modules. Journal of Number Theory, 26(3):308–324, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900874.

deWeger:1987:SED

[571] B. M. M. de Weger. Solving exponential diophantine equations using lattice Basis Reduction Algorithms. Journal of Number Theory, 26(3):325– 367, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X87900886.

Anonymous:1987:AIVb

[572] Anonymous. Author index for volume 26. Journal of Number Theory, 26 (3):368, July 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X87900898.

Anonymous:1987:PS

[573] Anonymous. Pages 1–112 (September 1987). Journal of Number Theory, 27(1):??, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1987:EBc

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900424.

Metsankyla:1987:SME

[575] Tauno Metsänkylä. A simple method for estimating the Iwasawa λinvariant. Journal of Number Theory, 27(1):1–6, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900436.

Nash:1987:FTA

[576] John C. M. Nash. Freiman’s theorem answers a question of Erdős. Journal of Number Theory, 27(1):7–8, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900448.

Cochrane:1987:TIV

[577] Todd Cochrane. On a trigonometric inequality of Vinogradov. Journal of Number Theory, 27(1):9–16, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8790045X.

Ito:1987:PED

[578] Hiroshi Ito. On a property of elliptic Dedekind sums. Journal of Number Theory, 27(1):17–21, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X87900461.

Filaseta:1987:IAA

[579] Michael Filaseta. The irreducibility of almost all Bessel polynomials. Journal of Number Theory, 27(1):22–32, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900473.

Alexandru:1987:CG

[580] Zaharescu Alexandru. On a conjecture of Graham. Journal of Number Theory, 27(1):33–40, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X87900485.

Poulakis:1987:ESC

[581] Dimitrios Poulakis. Évaluation d’une somme cubique de caractères. (French) [Evaluation of a cubic sum of characters]. Journal of Number Theory, 27(1):41–45, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900497.

Peng:1987:ATEa

[582] Chuang Peng. Addition theorems in elementary Abelian groups, I. Journal of Number Theory, 27(1):46–57, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900503.

Peng:1987:ATEb

[583] Chuang Peng. Addition theorems in elementary Abelian groups, II. Journal of Number Theory, 27(1):58–62, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X87900515.

Granville:1987:SGT

[584] Andrew Granville. Sophie Germain’s theorem for prime pairs p, 6p + 1. Journal of Number Theory, 27(1):63–72, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900527.

Ivic:1987:GDP

[585] Aleksandar Ivić. The general divisor problem. Journal of Number Theory, 27(1):73–91, September 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900539.

Buchmann:1987:ATL

[586] Johannes Buchmann and Jonathan W. Sands. An algorithm for testing Leopoldt’s conjecture. Journal of Number Theory, 27(1):92–105, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900540.

Moroz:1987:NIP

[587] B. Z. Moroz. On the number of integral points on a norm-form variety in a cube-like domain. Journal of Number Theory, 27(1):106–110, September 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X87900552.

Bhaskaran:1987:LE

[588] M. Bhaskaran. Letter to the editor. Journal of Number Theory, 27(1): 111–112, September 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900564.

Anonymous:1987:PO

[589] Anonymous. Pages 113–221 (October 1987). Journal of Number Theory, 27(2):??, October 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Gurak:1987:RGE

[590] S. Gurak. Rational and genus equivalence of forms. Journal of Number Theory, 27(2):113–143, October 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X87900576.

Mesirov:1987:CFE

[591] Jill P. Mesirov and Melvin M. Sweet. Continued fraction expansions of rational expressions with irreducible denominators in characteristic 24. Journal of Number Theory, 27(2):144–148, October 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900588.

Goldston:1987:FTR

[592] D. A. Goldston. On the function S(T) in the theory of the Riemann zeta-function. Journal of Number Theory, 27(2):149–177, October 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8790059X.

Hardy:1987:CMC

[593] Kenneth Hardy and Kenneth S. Williams. Congruences modulo 16 for the class numbers of complex quadratic fields. Journal of Number Theory, 27 (2):178–195, October 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900606.

Alon:1987:SS

[594] N. Alon. Subset sums. Journal of Number Theory, 27(2):196–205, October 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900618.

Zhenfu:1987:EZP

[595] Cao Zhenfu, Liu Rui, and Zhang Liangrui. On the equation ∑_{j=1}^{s}(1x_{j}) + (1(x_{1} … x_{s})) = 1 and Znám’s problem. Journal of Number Theory, 27 (2):206–211, October 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X8790062X.

Woodcock:1987:TVR

[596] C. F. Woodcock. A two variable Riemann zeta function. Journal of Number Theory, 27(2):212–221, October 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X87900631.

Anonymous:1987:PN

[597] Anonymous. Pages 223–353 (November 1987). Journal of Number Theory, 27(3):??, November 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Styer:1987:ESC

[598] Robert Styer. Eisenstein series over complex quadratic fields when the class number exceeds one. Journal of Number Theory, 27(3):223–237, November 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X87900643.

Horie:1987:III

[599] Kuniaki Horie. On Iwasawa λ^{−}-invariants of imaginary abelian fields. Journal of Number Theory, 27(3):238–252, November 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900655.

Lehman:1987:RPE

[600] J. Larry Lehman. Rational points on elliptic curves with complex multiplication by the ring of integers in Q(\sqrt{−7}). Journal of Number Theory, 27 (3):253–272, November 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900667.

Arenas:1987:ABS

[601] A. Arenas and P. Bayer. Arithmetic behaviour of the sums of three squares. Journal of Number Theory, 27(3):273–284, November 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900679.

Cogdell:1987:BC

[602] J. W. Cogdell and I. I. Piatetski-Shapiro. Base change for ≈ SL_{2}. Journal of Number Theory, 27(3):285–303, November 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X87900680.

Perlstadt:1987:SRS

[603] Marci A. Perlstadt. Some recurrences for sums of powers of binomial coefficients. Journal of Number Theory, 27(3):304–309, November 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X87900692.

Bedoya:1987:ACD

[604] Hernando Bedoya and Karl Otto Stöhr. An algorithm to calculate discrete invariants of singular primes in function fields. Journal of Number Theory, 27(3):310–323, November 1987. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X87900709.

Shorey:1987:PPR

[605] T. N. Shorey and C. L. Stewart. Pure powers in recurrence sequences and some related diophantine equations. Journal of Number Theory, 27 (3):324–352, November 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900710.

Anonymous:1987:AIVc

[606] Anonymous. Author index for volume 27. Journal of Number Theory, 27(3):353, November 1987. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X87900722.

Anonymous:1988:PJa

[607] Anonymous. Pages 1–118 (January 1988). Journal of Number Theory, 28 (1):??, January 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1988:EBa

[608] Anonymous. Editorial Board. Journal of Number Theory, 28(1):i, January 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8890114X.

Erdogdu:1988:CPR

[609] V. Erdoğdu. Coprimely packed rings. Journal of Number Theory, 28(1): 1–5, January 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X88901151.

Stratigopoulos:1988:PRG

[610] Dimitri Stratigopoulos. A property of ramification groups and numbers of algebraic finite extensions of valued fields. Journal of Number Theory, 28(1):6–16, January 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88901163.

Krasner:1988:LDA

[611] Marc Krasner. Local differents of algebraic and finite extensions of valued fields. Journal of Number Theory, 28(1):17–61, January 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88901175.

Moreno:1988:DIC

[612] Oscar Moreno. Discriminants and the irreducibility of a class of polynomials in a finite field of arbitrary characteristic. Journal of Number Theory, 28(1):62–65, January 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88901187.

Vaserstein:1988:EIS

[613] L. N. Vaserstein. Every integer is a sum or difference of 28 integral eighth powers. Journal of Number Theory, 28(1):66–68, January 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88901199.

Parimala:1988:WGC

[614] R. Parimala. Witt groups of conics, elliptic, and hyperelliptic curves. Journal of Number Theory, 28(1):69–93, January 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88901205.

Goldston:1988:SPI

[615] D. A. Goldston and Kevin S. McCurley. Sieving the positive integers by large primes. Journal of Number Theory, 28(1):94–115, January 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88901217.

Tong:1988:STA

[616] Jingcheng Tong. Segre’s theorem on asymmetric diophantine approximation. Journal of Number Theory, 28(1):116–118, January 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88901229.

Anonymous:1988:PF

[617] Anonymous. Pages 119–232 (February 1988). Journal of Number Theory, 28(2):??, February 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Ozeki:1988:EEU

[618] Michio Ozeki. Examples of even unimodular extremal lattices of rank 40 and their Siegel theta-series of degree 2. Journal of Number Theory, 28 (2):119–131, February 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88900595.

Loper:1988:RCD

[619] Alan Loper. On rings without a certain divisibility property. Journal of Number Theory, 28(2):132–144, February 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900601.

Sueyoshi:1988:CNF

[620] Yutaka Sueyoshi. A characterization of number fields by p-closed extensions. Journal of Number Theory, 28(2):145–151, February 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900613.

Cornell:1988:NSH

[621] Gary Cornell and Michael Rosen. A note on the splitting of the Hilbert class field. Journal of Number Theory, 28(2):152–158, February 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88900625.

Erdos:1988:SCI

[622] Paul Erdős, Melvyn B. Nathanson, and András Sárközy. Sumsets containing infinite arithmetic progressions. Journal of Number Theory, 28 (2):159–166, February 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88900637.

Gomez-Calderon:1988:PSV

[623] Javier Gomez-Calderon and Daniel J. Madden. Polynomials with small value set over finite fields. Journal of Number Theory, 28(2):167–188, February 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X88900649.

Sodaigui:1988:SGR

[624] Bouchaïb Sodaïgui. Structure galoisienne relative des anneaux d’entiers. Journal of Number Theory, 28(2):189–204, February 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900650.

Jia:1988:EOS

[625] Xing-De Jia. Exact order of subsets of asymptotic bases in additive number theory. Journal of Number Theory, 28(2):205–218, February 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88900662.

Stiller:1988:CAF

[626] P. F. Stiller. Classical automorphic forms and hypergeometric functions. Journal of Number Theory, 28(2):219–232, February 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900674.

Anonymous:1988:PMa

[627] Anonymous. Pages 233–370 (March 1988). Journal of Number Theory, 28(3):??, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Myerson:1988:DSU

[628] Gerald Myerson. Dedekind sums and uniform distribution. Journal of Number Theory, 28(3):233–239, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8890039X.

Hafner:1988:NOR

[629] James Lee Hafner. New omega results in a weighted divisor problem. Journal of Number Theory, 28(3):240–257, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900406.

Yokota:1988:DEF

[630] Hisashi Yokota. Denominators of Egyptian fractions. Journal of Number Theory, 28(3):258–271, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X88900418.

Yokota:1988:LDE

[631] Hisashi Yokota. Length and denominators of Egyptian fractions II. Journal of Number Theory, 28(3):272–282, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8890042X.

Kaminski:1988:CPU

[632] Michael Kaminski. Cyclotomic polynomials and units in cyclotomic number fields. Journal of Number Theory, 28(3):283–287, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88900431.

Bremner:1988:PBC

[633] Andrew Bremner. On power bases in cyclotomic number fields. Journal of Number Theory, 28(3):288–298, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900443.

Agarwal:1988:RRI

[634] A. K. Agarwal. Rogers–Ramanujan identities for n-color partitions. Journal of Number Theory, 28(3):299–305, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900455.

Diamond:1988:CSD

[635] H. Diamond, H. Halberstam, and H.-E. Richert. Combinatorial sieves of dimension exceeding one. Journal of Number Theory, 28(3):306–346, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88900467.

Chellali:1988:ACN

[636] Mustapha Chellali. Accélération de calcul de nombres de Bernoulli. (French) [Bernoulli number calculation acceleration]. Journal of Number Theory, 28(3):347–362, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X88900479.

Erdos:1988:NAG

[637] Paul Erdős and Michael E. Mays. On nilpotent but not abelian groups and abelian but not cyclic groups. Journal of Number Theory, 28(3):363– 368, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X88900480.

Anonymous:1988:AIVa

[638] Anonymous. Author index for volume 28. Journal of Number Theory, 28(3):369–370, March 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88900492.

Anonymous:1988:PMb

[639] Anonymous. Pages 1–99 (May 1988). Journal of Number Theory, 29(1): ??, May 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1988:EBb

[640] Anonymous. Editorial Board. Journal of Number Theory, 29(1):i, May 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8890087X.

Erdos:1988:PBD

[641] Paul Erdős and Melvyn B. Nathanson. Partitions of bases into disjoint unions of bases. Journal of Number Theory, 29(1):1–9, May 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900881.

Balasubramanian:1988:ACS

[642] R. Balasubramanian. On the additive completion of squares. Journal of Number Theory, 29(1):10–12, May 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X88900893.

Scott:1988:MMT

[643] Paul R. Scott. Modifying Minkowski’s theorem. Journal of Number Theory, 29(1):13–20, May 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X8890090X.

Dummit:1988:AEG

[644] D. S. Dummit and H. Kisilevsky. Abelian extensions generated by division points. Journal of Number Theory, 29(1):21–30, May 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900911.

Low:1988:SDC

[645] Lewis Low, Jane Pitman, and Alison Wolff. Simultaneous diagonal congruences. Journal of Number Theory, 29(1):31–59, May 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900923.

Iwaniec:1988:DCP

[646] H. Iwaniec and C. J. Mozzochi. On the divisor and circle problems. Journal of Number Theory, 29(1):60–93, May 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X88900935.

Elliott:1988:DNP

[647] P. D. T. A. Elliott and A. Sárközy. The distribution of the number of prime divisors of sums a + b. Journal of Number Theory, 29(1):94–99, May 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88900947.

Anonymous:1988:PJb

[648] Anonymous. Pages 101–233 (June 1988). Journal of Number Theory, 29 (2):??, June 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Wallace:1988:MPT

[649] D. I. Wallace. Maximal parabolic terms in the Selberg trace formula for SL(3, Z) \ SL(3, R)/SO(3, R). Journal of Number Theory, 29(2):101– 117, June 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X88900959.

Ramanathan:1988:GST

[650] K. G. Ramanathan. Generalisations of some theorems of Ramanujan. Journal of Number Theory, 29(2):118–137, June 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900960.

Aggarwal:1988:PVI

[651] Satish K. Aggarwal and D. P. Gupta. Positive values of inhomogeneous quadratic forms of signature −2. Journal of Number Theory, 29(2):138– 165, June 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X88900972.

Stroeker:1988:ASA

[652] R. J. Stroeker and N. Tzanakis. On the application of Skolem’s p-adic method to the solution of Thue equations. Journal of Number Theory, 29(2):166–195, June 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88900984.

Kalpazidou:1988:KCL

[653] Sofia Kalpazidou. Khintchine’s constant for Lüroth representation. Journal of Number Theory, 29(2):196–205, June 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900996.

Jean:1988:NTP

[654] Roger V. Jean. Number-theoretic properties of two-dimensional lattices. Journal of Number Theory, 29(2):206–223, June 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8890100X.

Stopple:1988:RLP

[655] Jeffrey Stopple. A reciprocity law for prime geodesics. Journal of Number Theory, 29(2):224–230, June 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X88901011.

Gerstein:1988:NSQ

[656] Larry J. Gerstein. A note on splitting quadratic forms. Journal of Number Theory, 29(2):231–233, June 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X88901023.

Anonymous:1988:PJc

[657] Anonymous. Pages 235–377 (July 1988). Journal of Number Theory, 29 (3):??, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Mauduit:1988:ENS

[658] Christian Mauduit. Sur l’ensemble normal des substitutions de longueur quelconque. (French) [On the normal set of substitutions of any length]. Journal of Number Theory, 29(3):235–250, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88901035.

Ribenboim:1988:REC

[659] P. Ribenboim. Remarks on exponential congruences and powerful numbers. Journal of Number Theory, 29(3):251–263, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88901047.

Burke:1988:SPS

[660] John R. Burke and William A. Webb. Sums of perturbed sequences of integers. Journal of Number Theory, 29(3):264–270, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88901059.

Hildebrand:1988:LVC

[661] Adolf Hildebrand. Large values of character sums. Journal of Number Theory, 29(3):271–296, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X88901060.

Thaine:1988:FLT

[662] Francisco Thaine. On Fermat’s last theorem and the arithmetic of Z[ζ_{p} + ζ_{p}^{−1}]. Journal of Number Theory, 29(3):297–299, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88901072.

Coster:1988:GCG

[663] Matthijs Coster. Generalisation of a congruence of Gauss. Journal of Number Theory, 29(3):300–310, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X88901084.

Hajela:1988:SRS

[664] D. Hajela. Some remarks on B_{h}[g] sequences. Journal of Number Theory, 29(3):311–323, July 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88901096.

Yuan:1988:DIF

[665] Wang Yuan. Diophantine inequalities for forms in an algebraic number field. Journal of Number Theory, 29(3):324–344, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88901102.

Effinger:1988:GPT

[666] Gove Effinger. A Goldbach 3-primes theorem for polynomials of low degree over finite fields of characteristic 2. Journal of Number Theory, 29(3):345– 363, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X88901114.

Harman:1988:MDA

[667] Glyn Harman. Metric diophantine approximation with two restricted variables. III. Two prime numbers. Journal of Number Theory, 29(3):364–375, July 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88901126.

Anonymous:1988:AIVb

[668] Anonymous. Author index for volume 29. Journal of Number Theory, 29(3):376–377, July 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88901138.

Anonymous:1988:PS

[669] Anonymous. Pages 1–107 (September 1988). Journal of Number Theory, 30(1):??, September 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1988:EBc

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88900212.

Bond:1988:UAE

[671] Robert J. Bond. Unramified abelian extensions of number fields. Journal of Number Theory, 30(1):1–10, September 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900224.

Gekeler:1988:PSP

[672] Ernst-Ulrich Gekeler. On power sums of polynomials over finite fields. Journal of Number Theory, 30(1):11–26, September 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900236.

Dafni:1988:MZQ

[673] Galia Dafni and Leonid Vaserstein. Matrix zeros of quadratic equations. Journal of Number Theory, 30(1):27–50, September 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900248.

Niederreiter:1988:LDL

[674] Harald Niederreiter. Low-discrepancy and low-dispersion sequences. Journal of Number Theory, 30(1):51–70, September 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X8890025X.

Petermann:1988:ATP

[675] Y.-F. S. Pétermann. About a theorem of Paolo Codecà’s and Ω-estimates for arithmetical convolutions. Journal of Number Theory, 30(1):71–85, September 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X88900261.

Mohanty:1988:IP

[676] S. P. Mohanty. Integer points of y^{2} = x^{3} − 4x + 1. Journal of Number Theory, 30(1):86–93, September 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X88900273.

Proinov:1988:DVC

[677] Petko D. Proinov and Vasil S. Grozdanov. On the diaphony of the Van der Corput–Halton sequence. Journal of Number Theory, 30(1):94–104, September 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X88900285.

Waterhouse:1988:TNG

[678] William C. Waterhouse. A tiny note on Gauss’s Lemma. Journal of Number Theory, 30(1):105–107, September 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900297.

Anonymous:1988:PO

[679] Anonymous. Pages 109–241 (October 1988). Journal of Number Theory, 30(2):??, October 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Zhang:1988:NTS

[680] Wen-Bin Zhang. On a number theoretic series. Journal of Number Theory, 30(2):109–119, October 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X8890011X.

Zhang:1988:DDB

[681] Wen-Bin Zhang. Density and O-density of Beurling generalized integers. Journal of Number Theory, 30(2):120–139, October 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900121.

Cougnard:1988:GAE

[682] Jean Cougnard. Génération de l’anneau des entiers des corps de classes de Q(i) de rayon impair et points de division de y^{2} = x^{2} − x. (French) [Generation of the ring of the integers of the field of classes of Q(i) of odd radius and points of division of y^{2} = x^{2} − x]. Journal of Number Theory, 30(2):140–155, October 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88900133.

Mignotte:1988:IAI

[683] Maurice Mignotte. An inequality about irreducible factors of integer polynomials. Journal of Number Theory, 30(2):156–166, October 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88900145.

Louboutin:1988:CFR

[684] Stéphane Louboutin. Continued fractions and real quadratic fields. Journal of Number Theory, 30(2):167–176, October 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900157.

Krieg:1988:ESF

[685] Aloys Krieg. Eisenstein-series on the four-dimensional hyperbolic space. Journal of Number Theory, 30(2):177–197, October 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900169.

Yokota:1988:PBE

[686] Hisashi Yokota. On a problem of Bleicher and Erdős. Journal of Number Theory, 30(2):198–207, October 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X88900170.

Filaseta:1988:EAS

[687] Michael Filaseta. An elementary approach to short interval results for k-free numbers. Journal of Number Theory, 30(2):208–225, October 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88900182.

Silverman:1988:WCA

[688] Joseph H. Silverman. Wieferich’s criterion and the abc-conjecture. Journal of Number Theory, 30(2):226–237, October 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900194.

Granville:1988:SRR

[689] Vincent Granville and Jean Paul Rasson. A strange recursive relation. Journal of Number Theory, 30(2):238–241, October 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900200.

Anonymous:1988:PN

[690] Anonymous. Pages 243–383 (November 1988). Journal of Number Theory, 30(3):??, November 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Shastri:1988:WGA

[691] Parvati Shastri. Witt groups of algebraic integers. Journal of Number Theory, 30(3):243–266, November 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X88900017.

Roux:1988:CCG

[692] R. Paysant-Le Roux. Calibre d’un corps global. (French) [Caliber of a global body]. Journal of Number Theory, 30(3):267–287, November 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X88900029.

Obata:1988:DNN

[693] Nobuaki Obata. Density of natural numbers and the Lévy group. Journal of Number Theory, 30(3):288–297, November 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900030.

Cogdell:1988:BCS

[694] J. W. Cogdell and I. I. Piatetski-Shapiro. Base change for the Saito– Kurokawa representations of PGSp(4). Journal of Number Theory, 30(3): 298–320, November 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88900042.

Denert:1988:OQA

[695] Marleen Denert and Jan Van Geel. Orders in quaternion algebras over global function fields having the cancellation property. Journal of Number Theory, 30(3):321–333, November 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X88900054.

Chou:1988:VSD

[696] Wun-Seng Chou, Javier Gomez-Calderon, and Gary L. Mullen. Value sets of Dickson polynomials over finite fields. Journal of Number Theory, 30 (3):334–344, November 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88900066.

Garcia:1988:FCF

[697] A. Garcia and J. F. Voloch. Fermat curves over finite fields. Journal of Number Theory, 30(3):345–356, November 1988. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X88900078.

Hemard:1988:MGT

[698] Denis Hemard. Modules galoisiens de torsion et plongements dans les Z_{p}-extensions. (French) [Galois torsion and embedding modules in Z_{p}-extensions]. Journal of Number Theory, 30(3):357–374, November 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8890008X.

Hoechsmann:1988:URA

[699] K. Hoechsmann and S. K. Sehgal. Units in regular abelian p-group rings. Journal of Number Theory, 30(3):375–381, November 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X88900091.

Anonymous:1988:AIVc

[700] Anonymous. Author index for volume 30. Journal of Number Theory, 30 (3):382–383, November 1988. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88900108.

Anonymous:1989:PJa

[701] Anonymous. Pages 1–89 (January 1989). Journal of Number Theory, 31 (1):??, January 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1989:EBa

[702] Anonymous. Editorial Board. Journal of Number Theory, 31(1):i, January 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900486.

Diaz:1989:GDT

[703] Guy Diaz. Grands degrés de transcendance pour des familles d’exponentielles. (French) [Great degrees of transcendence for families of exponentials]. Journal of Number Theory, 31(1):1–23, January 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900498.

Laurent:1989:EEP

[704] Michel Laurent. Équations exponentielles-polynômes et suites récurrentes linéaires, II. (French) [Exponential-polynomial equations and linear recurrent sequences, II]. Journal of Number Theory, 31(1):24–53, January 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900504.

Lecacheux:1989:UFC

[705] Odile Lecacheux. Unités d’une famille de corps cycliques réels de degré 6 liés à la courbe modulaire X_{1}(13). (French) [Units of a family of real cyclic bodies of degree 6 linked to the modular curve X_{1}(13)]. Journal of Number Theory, 31(1):54–63, January 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X89900516.

Kwong:1989:PCI

[706] Y. H. Harris Kwong. Periodicities of a class of infinite integer sequences modulo M. Journal of Number Theory, 31(1):64–79, January 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900528.

Cobeli:1989:GCU

[707] Cristian Cobeli, Marian Vâjâitu, and Alexandru Zaharescu. Grahm’s conjecture under Riemann hypothesis. Journal of Number Theory, 31(1):80– 87, January 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X8990053X.

Anonymous:1989:E

[708] Anonymous. Erratum. Journal of Number Theory, 31(1):88–89, January 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900541.

Anonymous:1989:PF

[709] Anonymous. Pages 91–241 (February 1989). Journal of Number Theory, 31(2):??, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

McDaniel:1989:DDS

[710] Wayne L. McDaniel. Difference of the digital sums of an integer base b and its prime factors. Journal of Number Theory, 31(2):91–98, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900139.

Tzanakis:1989:PST

[711] N. Tzanakis and B. M. M. de Weger. On the practical solution of the Thue equation. Journal of Number Theory, 31(2):99–132, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900140.

Livne:1989:CGR

[712] Ron Livné. On the conductors of mod /Galois representations coming from modular forms. Journal of Number Theory, 31(2):133–141, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900152.

Nakamula:1989:EUC

[713] Ken Nakamula. Elliptic units and the class numbers of non–Galois fields. Journal of Number Theory, 31(2):142–166, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900164.

Uehara:1989:RNI

[714] Tsuyoshi Uehara. On the 4-rank of the narrow ideal class group of a quadratic field. Journal of Number Theory, 31(2):167–173, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900176.

Nowak:1989:DPA

[715] Werner Georg Nowak. On a divisor problem in arithmetic progressions. Journal of Number Theory, 31(2):174–182, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900188.

Alladi:1989:MFS

[716] K. Alladi, P. Erdős, and J. D. Vaaler. Multiplicative functions and small divisors, II. Journal of Number Theory, 31(2):183–190, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8990019X.

Miyake:1989:ICE

[717] K. Miyake and H. Opolka. Indices of central embedding problems and applications. Journal of Number Theory, 31(2):191–200, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900206.

Gow:1989:SGL

[718] R. Gow. Some generalized Laguerre polynomials whose Galois groups are the alternating groups. Journal of Number Theory, 31(2):201–207, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X89900218.

Borel:1989:SSM

[719] Jean-Pierre Borel. Self-similar measures and sequences. Journal of Number Theory, 31(2):208–241, February 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X8990022X.

Anonymous:1989:PMa

[720] Anonymous. Pages 243–375 (March 1989). Journal of Number Theory, 31(3):??, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Daboussi:1989:PNT

[721] H. Daboussi. On the prime number theorem for arithmetic progressions. Journal of Number Theory, 31(3):243–254, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900711.

Kiss:1989:AF

[722] Péter Kiss and Ferenc Mátyás. An asymptotic formula for π. Journal of Number Theory, 31(3):255–259, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900723.

Schmutz:1989:PCE

[723] Eric Schmutz. Proof of a conjecture of Erdős and Turán. Journal of Number Theory, 31(3):260–271, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X89900735.

Yuan:1989:SZQ

[724] Wang Yuan. On small zeros of quadratic forms over finite fields. Journal of Number Theory, 31(3):272–284, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900747.

Erdos:1989:FAL

[725] P. Erdős and G. Tenenbaum. Sur les fonctions arithmétiques liées aux diviseurs consécutifs. (French) [On the arithmetic functions related to consecutive dividers]. Journal of Number Theory, 31(3):285–311, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900759.

Ichimura:1989:UPS

[726] Humio Ichimura and Masanobu Kaneko. On the universal power series for Jacobi sums and the vandiver conjecture. Journal of Number Theory, 31(3):312–334, March 1989. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X89900760.

Lemke:1989:ATI

[727] Paul Lemke and Daniel Kleitman. An addition theorem on the integers modulo n. Journal of Number Theory, 31(3):335–345, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900772.

Satoh:1989:ARF

[728] Junya Satoh. q-Analogue of Riemann’s ζ-function and q-Euler numbers. Journal of Number Theory, 31(3):346–362, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900784.

Marcu:1989:AE

[729] Danu Marcu. An arithmetic equation. Journal of Number Theory, 31 (3):363–366, March 1989. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X89900796.

Larcher:1989:DMT

[730] Gerhard Larcher. On the distribution of the multiples of an s-tupel of real numbers. Journal of Number Theory, 31(3):367–372, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900802.

Bremner:1989:LE

[731] A. Bremner. Letter to the editor. Journal of Number Theory, 31(3): 373, March 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X89900814.

Anonymous:1989:AIVa

[732] Anonymous. Author index for volume 31. Journal of Number Theory, 31(3):374–375, March 1989. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X89900826.

Anonymous:1989:PMb

[733] Anonymous. Pages 1–130 (May 1989). Journal of Number Theory, 32 (1):??, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1989:EBb

[734] Anonymous. Editorial Board. Journal of Number Theory, 32(1):i, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900930.

Wallace:1989:MPT

[735] D. I. Wallace. Minimal parabolic terms in the Selberg trace formula for SL(3, Z) \ SL(3, R)/SO(3, R). Journal of Number Theory, 32(1):1– 13, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X89900942.

Berge:1989:PDL

[736] Anne-Marie Bergé and Jacques Martinet. Sur un problème de dualité lié aux sphères en géométrie des nombres. (French) [On a duality problem related to spheres in number geometry]. Journal of Number Theory, 32(1): 14–42, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X89900954.

Almkvist:1989:PCA

[737] Gert Almkvist. Proof of a conjecture about unimodal polynomials. Journal of Number Theory, 32(1):43–57, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900966.

Cipra:1989:SLA

[738] Barry A. Cipra. On the Shimura lift, après Selberg. Journal of Number Theory, 32(1):58–64, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X89900978.

Kraft:1989:IIC

[739] James S. Kraft. Iwasawa invariants of CM fields. Journal of Number Theory, 32(1):65–77, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X8990098X.

Saias:1989:NES

[740] Eric Saias. Sur le nombre des entiers sans grand facteur premier. (French) [On the number of integers without large prime factor]. Journal of Number Theory, 32(1):78–99, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X89900991.

Girstmair:1989:IFR

[741] Kurt Girstmair. An index formula for the relative class number of an abelian number field. Journal of Number Theory, 32(1):100–110, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89901005.

Shan:1989:GTM

[742] Zun Shan and Edward T. H. Wang. Generalization of a theorem of Mahler. Journal of Number Theory, 32(1):111–113, May 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89901017.

Sarkozy:1989:FAT

[743] A. Sárközy. Finite addition theorems, I. Journal of Number Theory, 32(1):114–130, May 1989. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X89901029.

Anonymous:1989:PJb

[744] Anonymous. Pages 131–262 (June 1989). Journal of Number Theory, 32 (2):??, June 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Kersten:1989:ECM

[745] Ina Kersten and Johannes Michaliček. Z_{p}-extensions of complex multiplication fields. Journal of Number Theory, 32(2):131–150, June 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900231.

Hafner:1989:MSR

[746] James Lee Hafner and Aleksandar Ivić. On the mean-square of the Riemann zeta-function on the critical line. Journal of Number Theory, 32 (2):151–191, June 1989. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X89900243.

Mai:1989:LBR

[747] Liem Mai. Lower bounds for the ranks of CM types. Journal of Number Theory, 32(2):192–202, June 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X89900255.

Stern:1989:NGG

[748] Leonid Stern. On the norm groups of global fields. Journal of Number Theory, 32(2):203–219, June 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X89900267.

Clark:1989:PCA

[749] W. Edwin Clark and Larry W. Lewis. Prime cyclic arithmetic codes and the distribution of power residues. Journal of Number Theory, 32(2):220– 225, June 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X89900279.

Kurihara:1989:AEU

[750] Fumika Kurihara. On the p-adic expansion of units of cyclotomic fields. Journal of Number Theory, 32(2):226–253, June 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900280.

Spyropoulos:1989:EEN

[751] Konstantine Spyropoulos. Euler’s equation ϕ(x) = k with no solution. Journal of Number Theory, 32(2):254–256, June 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900292.

Turnwald:1989:NHP

[752] Gerhard Turnwald. On the nonvanishing of homogeneous product sums. Journal of Number Theory, 32(2):257–262, June 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900309.

Anonymous:1989:PJc

[753] Anonymous. Pages 263–387 (July 1989). Journal of Number Theory, 32 (3):??, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Bremner:1989:DLC

[754] Andrew Bremner and Richard K. Guy. The delta-lambda configurations in tiling the square. Journal of Number Theory, 32(3):263–280, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900838.

Indlekofer:1989:GMA

[755] Karl-Heinz Indlekofer and Imre Kátai. Generalized moments of additive functions. Journal of Number Theory, 32(3):281–288, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8990084X.

Fei:1989:SNL

[756] Xu Fei. Spinor norms of local integral rotations in characteristic 2. Journal of Number Theory, 32(3):289–296, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900851.

Knopfmacher:1989:SEA

[757] A. Knopfmacher and J. Knopfmacher. Series expansions in p-adic and other non–Archimedean fields. Journal of Number Theory, 32(3):297– 306, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X89900863.

Hinz:1989:SFV

[758] Jürgen G. Hinz. Square-free values of cubic polynomials in algebraic number fields. Journal of Number Theory, 32(3):307–320, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900875.

Nart:1989:BCC

[759] Enric Nart. The Bloch complex in codimension one and arithmetic duality. Journal of Number Theory, 32(3):321–331, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900887.

Xianke:1989:CMC

[760] Zhang Xianke. Congruences modulo 8 for class numbers of general quadratic fields Q(\sqrt{m}) and Q(\sqrt{−m}). Journal of Number Theory, 32 (3):332–338, July 1989. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X89900899.

Elliott:1989:VDD

[761] P. D. T. A. Elliott. The value distribution of differences of additive arithmetic functions. Journal of Number Theory, 32(3):339–370, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900905.

Kersten:1989:VCE

[762] Ina Kersten and Johannes Michaliček. On Vandiver’s conjecture and Z_{p}- extensions of Q(ζ_{p^{n}}). Journal of Number Theory, 32(3):371–386, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900917.

Anonymous:1989:AIVb

[763] Anonymous. Author index for volume 32. Journal of Number Theory, 32 (3):387, July 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X89900929.

Anonymous:1989:PS

[764] Anonymous. Pages 1–106 (September 1989). Journal of Number Theory, 33(1):??, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:1989:EBc

[765] Anonymous. Editorial Board. Journal of Number Theory, 33(1):i, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900553.

Yeung:1989:CBC

[766] Kit Ming Yeung. On congruences for binomial coefficients. Journal of Number Theory, 33(1):1–17, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900565.

Gerth:1989:CRQ

[767] Frank Gerth. The 4-class ranks of quadratic extensions of certain real quadratic fields. Journal of Number Theory, 33(1):18–31, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900577.

Hironaka:1989:LDA

[768] Yumiko Hironaka and Fumihiro Sato. Local densities of alternating forms. Journal of Number Theory, 33(1):32–52, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900589.

Leep:1989:LDA

[769] D. B. Leep, J.-P Tignol, and N. Vast. The level of division algebras over local and global fields. Journal of Number Theory, 33(1):53–70, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900590.

Ruzsa:1989:APP

[770] Imre Z. Ruzsa. An additive problem for powers of primes. Journal of Number Theory, 33(1):71–82, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X89900607.

Durant:1989:SNR

[771] William R. Bon Durant. Spinor norms of rotations of local integral quadratic forms. Journal of Number Theory, 33(1):83–94, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900619.

Gilmer:1989:SDI

[772] Robert Gilmer. Sets that determine integer-valued polynomials. Journal of Number Theory, 33(1):95–100, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900620.

Leu:1989:DCR

[773] Ming-Guang Leu. On a determination of certain real quadratic fields of class number two. Journal of Number Theory, 33(1):101–106, September 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900632.

Anonymous:1989:PO

[774] Anonymous. Pages 107–265 (October 1989). Journal of Number Theory, 33(2):??, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Styer:1989:HTA

[775] Robert Styer. Hecke theory over arbitrary number fields. Journal of Number Theory, 33(2):107–131, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X89900012.

Solinas:1989:JDD

[776] Jerome A. Solinas. On the joint distribution of digital sums. Journal of Number Theory, 33(2):132–151, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900024.

Varbanec:1989:DGI

[777] P. D. Varbanec and P. Zarzycki. Divisors of the Gaussian integers in an arithmetic progression. Journal of Number Theory, 33(2):152–169, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/article/
pii/0022314X89900036.

Geluk:1989:ATT

[778] J. L. Geluk. An Abel–Tauber theorem for partitions, II. Journal of Number Theory, 33(2):170–181, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0022314X89900048.

Hensley:1989:HDS

[779] Doug Hensley. The Hausdorff dimensions of some continued fraction Cantor sets. Journal of Number Theory, 33(2):182–198, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X8990005X.

Reznick:1989:SSP

[780] Bruce Reznick. The sum of the squares of the parts of a partition, and some related questions. Journal of Number Theory, 33(2):199–208, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900061.

Gekeler:1989:SNI

[781] Ernst-Ulrich Gekeler. Some new identities for Bernoulli–Carlitz numbers. Journal of Number Theory, 33(2):209–219, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900073.

Kolitsch:1989:RBC

[782] Louis Worthy Kolitsch. A relationship between certain colored generalized Frobenius partitions and ordinary partitions. Journal of Number Theory, 33(2):220–223, October 1989. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X89900085.

Estes:1989:FQO

[783] Dennis R. Estes and Gordon Nipp. Factorization in quaternion orders. Journal of Number Theory, 33(2):224–236, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900097.

Waterhouse:1989:DAC

[784] William C. Waterhouse. The determination of all Carlitz sets. Journal of Number Theory, 33(2):237–242, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900103.

Petho:1989:DER

[785] Attila Pethö and Robert R. F. Tichy. On digit expansions with respect to linear recurrences. Journal of Number Theory, 33(2):243–256, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900115.

Parish:1989:RTC

[786] James L. Parish. Rational torsion in complex-multiplication elliptic curves. Journal of Number Theory, 33(2):257–265, October 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X89900127.

Anonymous:1989:PN

[787] Anonymous. Pages 267–363 (November 1989). Journal of Number Theory, 33(3):??, November 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Sheingorn:1989:HRP

[788] Mark Sheingorn. Hyperbolic reflections on Pell’s equation. Journal of Number Theory, 33(3):267–285, November 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900644.

Cochrane:1989:SZQ

[789] Todd Cochrane. Small zeros of quadratic forms modulo p. Journal of Number Theory, 33(3):286–292, November 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022314X89900656.

DeKoninck:1989:AFL

[790] Jean-Marie DeKoninck, Imre Kátai, and Armel Mercier. Additive functions and the largest prime factor of integers. Journal of Number Theory, 33(3):293–310, November 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0022314X89900668.

Fouvry:1989:ESM

[791] Etienne Fouvry and Henryk Iwaniec. Exponential sums with monomials. Journal of Number Theory, 33(3):311–333, November 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0022314X8990067X.

Goh:1989:DRR

[792] William M. Y. Goh. Dynamical representation of real numbers and its universality. Journal of Number Theory, 33(3):334–355, November 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900681.

Gerardin:1989:IIQ

[793] Paul Gérardin and Chung Yuan Lin. An identity involving quadratic characters. Journal of Number Theory, 33(3):356–361, November 1989. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/pii/
0022314X89900693.

Anonymous:1989:AIVc

[794] Anonymous. Author index for volume 33. Journal of Number Theory, 33 (3):362–363, November 1989. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X8990070X.